### v1

# Criteria for optimal entanglement-assisted long baseline imaging protocols

Abstract
Quantum-entanglement-assisted telescopy protocol have been proposed and studied as a means to extend the baseline of optical interferometric telescopes. However, determining the optimality of specific quantum-assisted protocols remains an open question. We introduce a new framework for systematically characterizing different entanglement-assisted telescopy protocols by integrating tools from quantum metrology with the superselection rule (SSR) established in quantum information theory. We demonstrate that the estimation of interference fringes in a quantum telescopy can be objectively quantified and compared using the quantum Fisher information under the SSR constraint. Our framework demonstrates that interference fringe estimation in quantum telescopy can be objectively evaluated and compared using the quantum Fisher information under the SSR constraint. This approach defines the fundamental limits of astronomical estimation problems with finite entanglement resources. Building on this foundation, we propose new entanglement-assisted protocols that outperform previously known methods. Moreover, we show that these protocols are compatible with linear optical technology and could inspire the development of improved quantum telescopy schemes for near-term, lossy, repeater-less quantum networks.
preprint: APS/123-QED
I Introduction
The basic principles of stellar interferometry involve the coherent measurement of light using distinct collection telescopes to form an effective large-scale imaging system. This method enhances the angular resolution well beyond the diffraction limit of a single telescope, restricted by its aperture [1, 2]. Unlike interferometric imaging in the radio frequency spectrum, where Earth-sized telescope arrays have been hugely successful, e.g. imaging black holes [3, 4], optical systems face major scaling limitations. In the standard ”direct detection” approach, photons collected at two telescopes are coherently recombined for interference [5, 6]. However, this method requires a stable optical channel between the telescopes, and current technologies achieve a baseline of about 300 meters [5]. Furthermore, methods based on optical homodyning using local oscillators have fundamental difficulties in achieving high signal-to-noise ratios [7]. These challenges have motivated interferometric telescopy assisted by quantum network technologies [8, 9].
The first quantum-networking-enhanced optical interferometry method was proposed by Gottesman, Jennewein, and Croke (GJC) [10], where they show that high signal-to-noise ratio local measurements can be achieved at arbitrarily long baselines by utilizing single-photon-entangled (SPE) states provided by quantum repeater networks. This surprising result was rigorously explained by Tsang [11] using tools from quantum optics and quantum estimation theory, where they highlighted that the non-local (NL) property of the ancilla single-photon states is essential for the enhancement in quantum telescopy. By further exploring the non-local property of different ancilla states, additional entanglement-assisted protocols have been proposed such as using quantum memories and error correction to reduce the number of required single-photon entangled states [12, 13], implementing control gates to boost the success rate of the original GJC protocol [14], or using multiple copies of SPE states to asymptotically achieve optimal performance[15, 16]. Recently, continuous variable quantum teleportation using two-mode squeeze vacuum states has shown promising results [17, 18]. Although there are many protocols supported by entanglement, there is no explicit way to determine whether any of these protocols is optimal, especially regarding their resource efficiency. Here we provide a systematic study of entanglement-assisted protocols by firstly recognizing that one must consider both locality and superselection rule (SSR) as constraints in order to clearly quantify the performance of an entanglement-assisted interferometric telescopy protocol.
The locality constraint, as introduced in [11], is a well-recognized limitation in classical astronomical imaging operating without direct detection. Non-local resources from entangled-assisted state, however, enable overcoming this constraint. Nevertheless, the locality constraint alone is insufficient for comparing different quantum telescopy protocols. A previously overlooked yet critical constraint arises from the absence of a phase reference—formally described as the superselection rule (SSR)[19]. By examining the quantum Fisher information of astronomical photon states under the SSR constraint, we demonstrate that an SSR associated with $U(1)$ symmetry and photon number conservation[20, 21] imposes additional restrictions on the amount of information extractable from astronomical interference measurements. To address this limitation, we systematically analyze and utilize various ancilla states as a phase reference (PR) in order to effectively overcome the SSR constraint. Importantly, this framework allows us to explore the fundamental limits of ancilla-state-assisted schemes, quantify the effectiveness of different ancilla states, and propose improved entanglement-assisted protocols. Our study provides a deeper understanding of the limitations and potential of quantum telescopy protocols in general [15, 17]. Furthermore, we discuss the practical implementation of these protocols using linear optics and classical communication, offering new insights into the realization of quantum telescopy protocols in near-term, repeater-less quantum networks. This approach builds on established techniques from linear optical quantum computation (LOQC) [22]
II Preliminaries
II.1 Quantum estimation problem
The astronomical object sensed by two telescopes can be modeled as a bipartite quasi-monochromatic weak thermal source with the mean photon number $\epsilon\ll 1$ (typical in optical interferometry [23]), which can be represented as the density operator:
$\rho_{s}=(1-\epsilon)\rho_{s}^{(0)}+\epsilon\rho_{s}^{(1)}+O(\epsilon^{2})\rho_{s}^{(>1)},$ (1)
where the single-photon term $\rho_{s}^{(1)}$, as the leading order term encoded with the complexity visibility $g$ can be expressed in the basis $\{|0\rangle_{A}|1\rangle_{B},|1\rangle_{A}|0\rangle_{B}\}$ as:
$\rho_{s}^{(1)}=\frac{1}{2}\begin{pmatrix}1&g\\ g^{*}&1\end{pmatrix},$ (2)
with the two telescopes labeled by $A$ and $B$ (Fig. 1), comprising the optical interferometer. Since the intensity distribution of an astronomical source can be inferred from the complex visibility $g=|g|e^{i\theta}$ through the Van Cittert-Zernike theorem[2, 1], achieving a precise estimation of the complex visibility $g$ is the primary objective of an astronomical interferometric imaging problem [5].
Figure 1: (a): Direct detection: light collected at two telescopes A and B are combined and interfered with each other. (b) Quantum entanglement assisted detection: a source provides a shared quantum resource to each telescope which allows interometric imaging via local linear optical circuit $\mathcal{U}$ and local measurements.
To quantify the amount of information about complex visibility $g$ contained in the source state described in Eq. 2, we consider the quantum Fisher information (QFI) matrix[24] $\mathbb{H}_{\mu,\nu}$ of estimating parameter $\mu,\nu\in(|g|,\theta)$ given a source state $\rho_{s}$. The inverse of the QFI matrix serves as a lower bound on the covariant matrix $\text{Cov}_{\mu,\nu}$ of estimating parameter $\mu,\nu\in(|g|,\theta)$ according to Cramer-Rao bound[25]. Note that QFI is also an upper bound Fisher information (FI) matrix $\mathbb{F}$, which describes the amount of information obtained from specific arbitrary measurements. This relationship leads to the following chain of inequalities:
$\mathbb{F}\leq\mathbb{H}\leq 1/\text{Cov}$ (3)
with $\mathbb{F}$ representing the Fisher information matrix. In the subsequent discussion, we will simply use $\mathbb{F}_{\mu}$ and $\mathbb{H}_{\mu}$ to refer to the diagonal elements of this matrix.
Definition 1.
Optimal quantum telescopy: the quantum Fisher information (QFI) of estimating $|g|$ and $\theta$ in source state $\rho_{s}$ is given as:
$\mathbb{H}_{|g|}[\rho_{s}^{AB}]=\frac{1}{1-|g|^{2}}\epsilon+O(\epsilon^{2}),~{}\mathbb{H}_{\theta}[\rho_{s}^{AB}]=|g|^{2}\epsilon+O(\epsilon^{2}).$ (4)
This quantity is derived in the appendix A, assuming that no constraint is present and thus will serve as an upper bound for any quantum telescopy protocol.
Remark.
The above QFI is achievable in the ideal case, including direct interference with noise-less and loss-less optical channel, continuous variable teleportation[17] and the control-gate-based protocol [14] with perfect phase reference and quantum memory.
The optimal performance of quantum telescopy protocol discussed above is achievable either through an ideal direct interference scheme or by assuming the existence of a perfect phase reference, without addressing how such a reference can be realized. In the following, we formally define the estimation problem for quantum telescopy under the superselection rule (SSR) constraint and outline how to overcome this limitation using a phase reference in the most resource-efficient manner, which also allows us to quantify and compare different ancilla-assisted quantum telescopy protocol.
II.2 Superselection rule
Let us now specify the theoretical description of the quantum telescopy in the presence of the super-selection rule. In a quantum optics experiment, the quantum states of optical modes are always defined relative to a phase reference [19]. Consider a state $|\psi\rangle$ defined relative to Alice’s phase reference frame as $|\psi\rangle=\sum_{n=0}c_{n}|n\rangle$ in the photon number basis. However, the state will be described differently in Bob’s phase references by a unitary transformation $U(\phi)=e^{i\phi\hat{N}}$, where $\phi$ is the angle between Bob’s and Alice’s phase reference frames, and $\hat{N}$ is the number operator. More specifically, the state viewed by Bob becomes $|\psi^{\prime}\rangle=U(\phi)|\psi\rangle=\sum_{n=0}c_{n}e^{in\phi}|n\rangle$. The lack of phase information $\phi$ restricts the kind of state prepared by Alice that can be recognized by Bob, and as we detailed in the Appendix B, a quantum system discussed in the photon number basis can be conveniently described in a Hilbert space decomposed as $H=\sum_{n=0}^{\infty}P_{n}HP_{n}:=\oplus_{n=0}^{\infty}H_{n}$ [26], where $P_{n}=|n\rangle\langle n|$ is the projector onto the $n$-photon number basis.
Similarly, for a bipartite system of two specially separated parties (telescope A and B), if there is no explicit phase reference between them, the Hilbert space will take the decomposition[27, 28, 29]
$H=\bigoplus_{N=0}^{\infty}H_{N}=\bigoplus_{N=0}^{\infty}\bigoplus_{n=1}^{N}H^{A}_{n}\otimes H^{B}_{N-n}$ (5)
under the SSR constraint and effectively, a restriction will be imposed to the possible bipartite state in the system:
Lemma 1.
A bipartite state $\rho^{AB}$ under the photon-number superselection rule will be represented as[29]:
$\rho^{AB}\Rightarrow\mathcal{E}(\rho^{AB})=\sum_{n,m=0}^{\infty}(P^{A}_{n}\otimes P^{B}_{m})\rho^{AB}(P^{A}_{n}\otimes P^{B}_{m})$ (6)
where $P^{A(B)}_{n}=|n\rangle\langle n|$ represents projector on $H^{A}_{n}$.
III Quantum Fisher Information with superselection rule
As an important consequence of Lemma 10, with SSR, the first order term $\rho^{(1)}_{s}$ in Eq. 2 will be completely dephased and become independent of $g$ and
$\mathcal{E}(\rho_{s})=(1-\epsilon)\mathbb{I}^{(0)}+\epsilon\frac{\mathbb{I}^{(1)}}{2}+O(\epsilon^{2})\mathcal{E}(\rho_{s}^{(>1)})$ (7)
Therefore, the achievable QFI scales at most $\epsilon^{2}$, which could already explain why an intensity interferometer is not efficient in very-long baseline interferometry[30, 5], and extra resources is required if direct detection is not possible due to noise or loss in optical channel.
The extra resources required to overcome superselection rule has been known as the phase reference (PR) and has been extensively discussed in quantum information science [19]. Next, we will turn to analyze these particular phase reference resources that are needed to supply between telescopes A and B to “lift” the $U(1)-$ superselection rule, which could be in general be described by an additional ancilla state $\rho_{a}$ that acts as a phase reference for the source state $\rho_{s}$[19]. Analog to the nonlocal resources (NL) source used to overcome locality constrains, in the following, we will study the phase reference (PR) source to fight against the superselection rule constrains. Formally speaking, when an ancilla state is included, the composite system takes the following form under SSR constraint:
$\mathcal{E}(\rho^{AB})$ $=\mathcal{E}(\rho_{s}^{AB}\otimes\rho_{a}^{AB})$
$=\sum_{n,m=0}^{\infty}(P^{A}_{n}\otimes P^{B}_{m})\rho_{s}^{AB}\otimes\rho_{a}^{AB}(P^{A}_{n}\otimes P^{B}_{m})$ (8)
Since $\rho_{a}^{AB}$ has no dependence on $g$, the additivity and monotonicity of QFI under post-processing ensure that:
Lemma 2.
The QFI of the composite system in the presence of SSR is upper bounded by the optimal one
$\mathbb{H}_{|g|}[\mathcal{E}(\rho^{AB})]\leq\mathbb{H}_{|g|}[\rho_{s}^{AB}]\quad\mathbb{H}_{\theta}[\mathcal{E}(\rho^{AB})]\leq\mathbb{H}_{\theta}[\rho_{s}^{AB}]$ (9)
Moreover, in order to find the optimal ancilla state that can saturate or asymptotically saturate the above inequality, it is convenient to define the following QFI ratio to quantify the performance of different ancilla states.
Definition 2.
QFI ratio: We evaluate an ancilla-assisted quantum telescopy protocol by quantifying the ratio of its QFI to the optimal QFI as:
$\mathbb{h}[\rho^{AB}_{a}]=\min\{\frac{\mathbb{H}_{|g|}[\mathcal{E}(\rho^{AB})]}{\mathbb{H}_{|g|}[\rho_{s}^{AB}]},\frac{\mathbb{H}_{\theta}[\mathcal{E}(\rho^{AB})]}{\mathbb{H}_{\theta}[\rho_{s}^{AB}]}\}$ (10)
[…]
In this work, we describe the most general ancilla state using the following notation:
$|\psi\rangle_{a}=\sum_{n,m=0}f_{n,m}|n_{A},m_{B}\rangle,$ (11)
where we use the notation $|n_{A},m_{B}\rangle$ to represent a state with $n$ photons at $A$ and $m$ photons at $B$, without specifying the detailed mode configurations. This abstraction encompasses states for various cases, such as:
  1. 1.
    A state $|n\rangle_{A}|m\rangle_{B}$, where all $n$ photons at $A$ occupy a single mode, and all $m$ photons at $B$ occupy another single mode.
  2. 2.
    A state $|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes m}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes m}_{B}$ with $(n+m)$ mode at $A$ and $B$, where one photon in each of the fisrt $n$ mode at $A$ and in each of the last $m$ model at $B$;
  3. 3.
    A state with superposition of different mode configurations: $\sum_{\pi}\sqrt{p(\pi)}\pi(|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes m}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes m}_{B})$, where $\pi$ is a permutation over mode configuration.
The abstraction provided by this notation is particularly useful for discussing the quantum Fisher information ratio, as the mode configuration does influence the calculation, as detailed in later sections and Appendix C.
Proposition 1.
Given an ancilla state $\rho_{a}=|\psi\rangle\langle\psi|$ of the form in Eq. 8, the QFI in estimating $|g|$ and $\theta$ of source state $\rho_{s}$ in the presence of SSR is given as:
$\mathbb{H}_{|g|}[\mathcal{E}(\rho^{AB})]=\frac{\epsilon}{1-|g|^{2}}\sum_{n,m=1}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}$
$\mathbb{H}_{\theta}[\mathcal{E}(\rho^{AB})]=\epsilon|g|^{2}\sum_{n,m=1}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}$
$\mathbb{h}[\rho^{AB}_{a}]=\sum_{n,m=1}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}$ (12)
Here we neglect the higher term $O(\epsilon)$ for simplicity. So $\mathbb{h}=0$ should actually be understood as $\mathbb{h}=O(\epsilon)$.
To make sense of our discussion, some constraints on the shared ancilla state $|\psi\rangle_{a}$ need to be imposed, a natural one is to limit the maximum photon number to some fixed value $N$, with this we have
Theorem 1.
For ancilla state $|\psi\rangle_{a}$ using at most $N$-photon in total, we have:
$\mathbb{h}[\rho^{AB}_{a}]\leq\cos(\frac{\pi}{N+2})\stackrel{{\scriptstyle\text{Large $N$}}}{{\approx}}(1-\frac{\pi^{2}}{2N^{2}})$ (13)
Another way of imposing the constraint could be on the average photon number $\langle N\rangle=\sum_{n,m}|f_{n,m}|^{2}(n+m)$, where a similar bound can be proven.
Corollary 1.
For ancilla state $|\psi\rangle_{a}$ with average photon number $\langle N\rangle$, we have:
$\mathbb{h}[\rho^{AB}_{a}]\leq\cos(\frac{\pi}{\langle N\rangle+2})\stackrel{{\scriptstyle\text{Large $\langle N\rangle$}}}{{\approx}}(1-\frac{\pi^{2}}{2\langle N\rangle^{2}})$ (14)
We leave the derivation of these upper bounds in the appendix C but would like to point out that the optimal ancilla states that saturate the above bound have the form,
$|\psi\rangle_{a}=\sum_{n=0}^{N}f_{n}|n_{A},(N-n)_{B}\rangle,$ (15)
This indicates that the optimal ancilla state always has coefficients $f_{n,m}=\delta_{n,N-m}f_{n}$, where $N$ is the amount of resources used in the ancilla system. And the Quantum Fisher information ratio gets simplified to:
$\mathbb{h}[\rho^{AB}_{a}]=\sum_{n=1}^{N}\frac{2|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}$ (16)
Ancilla state: $|\psi\rangle_{a}$ QFI ratio: $\mathbb{h}$ Realization Resource
LOCC Linear optics NL  PR
GJC scheme[10] $\frac{1}{\sqrt{2}}(|0\rangle_{A}|1\rangle_{B}+|1\rangle_{A}|0\rangle_{B})$ $\frac{1}{2}$  ✓
$N$-copy SPE[16, 15] $\frac{1}{\sqrt{2^{N}}}(|0\rangle_{A}|1\rangle_{B}+|1\rangle_{A}|0\rangle_{B})^{\otimes N}$ $1-\frac{1}{N+1}$  ✓
KLM scheme $\frac{1}{\sqrt{N+1}}\sum\limits_{n=0}^{N}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$ $1-\frac{1}{N+1}$  ✓
Modified-KLM1 $\sum\limits_{n=0}^{N}f^{\text{Itri}}_{n}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$ $\approx 1-2\frac{\log(N)}{N^{2}}$  ✓
Modified-KLM2 $\sum\limits_{n=0}^{N}f^{\text{Atri}}_{n}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$ $\approx 1-12\frac{1}{N^{2}}$  ✓
Optimal-KLM $\sum\limits_{n=0}^{N}\frac{\sqrt{2}\sin(\frac{n+1}{N+2}\pi)}{\sqrt{N+2}}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$ $\approx 1-\pi^{2}\frac{1}{N^{2}}$  ✓
CV-teleportation[17, 18] $\sum\limits_{n=0}^{\infty}\frac{\sqrt{2\langle N\rangle^{n}}}{\sqrt{(2+\langle N\rangle)^{n+1}}}|n\rangle_{A}|n\rangle_{B}$ $\geq 1-\frac{1}{\epsilon\langle N\rangle}^{\dagger}$  ✗
Two-photon state $\frac{1}{\sqrt{2}}(|01\rangle_{A}|10\rangle_{B}+|10\rangle_{A}|01\rangle_{B})$ $\frac{1}{2}^{\ddagger}$  ✗
Coherent state $|\alpha\rangle_{A}|\alpha\rangle_{B}$ $\approx 1-\frac{1}{2|\alpha|^{2}}$  ✓
Table 1: Summary on the QFI ratio $\mathbb{h}$ computed for different ancilla states used in the quantum telescopy protocol. The realizations of the QFI-ratios of these schemes are categorized by whether they could be achieved with local operation with classical communication (LOCC), and linear optics. The ancilla state is considered as a non-local (NL) resource, phase-reference (PR) resource (to lift the SSR), or both. Importantly, not all non-local ancilla states can be used to build a phase reference and overcome the SSR, and there exist classically correlated states, that can be used to build a phase reference, but can not help to overcome locality constraints; Therefore, a good ancilla state should be both NL resource and PR resource. : assuming an extra perfect share-phase reference exists, i.e., $|\alpha\rangle_{A}|\alpha\rangle_{B}$ with $|\alpha|\rightarrow\infty$; : with preprocessing.
We now provide explicit examples of different ancilla-assisted quantum telescopy protocols and their corresponding quantum Fisher information under SSR constraints. It is worth noting that the direct interference scheme does not belong to this category, as it is not subject to SSR constraints. The performance of different ancilla states is summarized in Table 1, and the comparison to the numerical optimal ancilla state is given in Fig. 2.
Figure 2: The Quantum Fisher information (QFI) ratio as a function of photons number used in the ancilla state. The different entangled ancilla schemes are described in the main text.
IV Comparing various protocols
As mentioned in the preliminary section, in the absence of ancilla states, it is straightforward to verify that
$\mathbb{h}[\text{Null}]=0,$ (17)
This corresponds to the performance of intensity interferometry, where no ancilla states are used, and the “fringe” estimation is carried out after local intensity measurements. This observation motivates the exploration of ancilla states and their role in enhancing the performance of quantum telescopy protocols, which will be elaborated upon in the following subsections.
IV.1 GJC scheme
First we consider the original scheme proposed by Gottesman-Jennewein-Croke in [10].
With the single-particle entangled (SPE) ancilla state $|\psi\rangle_{a}=\frac{1}{\sqrt{2}}(|0\rangle_{A}|1\rangle_{B}+|1\rangle_{A}|0\rangle_{B})$, using Proposition 1, one could easily find:
$\mathbb{h}[|\psi\rangle_{a}]=\frac{1}{2}.$ (18)
Importantly, the quantum Fisher information (ratio) above is known to be achievable using correlated inference measurements (explicit protocols exist that achieve a Fisher information ratio saturating this bound). However, as noted in the original paper [10] and many subsequent discussions, such a scheme will succeed only $50\%$ of the time thereby providing a Fisher information ratio of $\frac{1}{2}$. Here we show that this $\frac{1}{2}$ ratio is not only any bound from a specific measurement protocol (the correlated interference measurement in [10] ), but also fundamental on the level of quantum fisher information when superselection rule is properly reconginized.
IV.2 N-copy SPE
A modification of the original GJC scheme uses an N-copy-version[15] of the SPE, where the ancilla state takes the form of $|\psi\rangle^{\otimes N}_{a}=\frac{1}{\sqrt{2^{N}}}(|0\rangle_{A}|1\rangle_{B}+|1\rangle_{A}|0\rangle_{B})^{\otimes N}$.
The discussion on N-copy-SPE state has been attempted in several works [16, 15], however, its performance has not been carefully studied yet  [^1] It was argued in [15], that the upper bound FI scales $1-\frac{1}{N+1}$, however, the conclusion is drawn assuming all photons enter the same telescopes, which could only give an loose bound scales $1-\frac{1}{2^{N}}$ Following Proposition 1, we can immediately show that the QFI-ratio using N-copy SPE is:
$|\psi\rangle^{\otimes N}_{a}=\sum_{n=0}^{N}\sqrt{\frac{{N\choose n}}{2^{N}}}|n_{A},(N-n)_{B}\rangle,$ (19)
where, we use $|n_{A},(N-n)_{B}\rangle$ to represent a state with $n$ photon at telescope $A$ and $m$ photon at telescope B for simplicity without specifying its mode configurations (i.e., it is indeed a superposition state of different mode configurations). With this, we can compute the QFI-ratio as:
$\mathbb{h}[|\psi\rangle^{\otimes N}_{a}]=\frac{N}{N+1}$
Agreeing with the Fisher information computed and conjectured in [16, 15]. However, this QFI cannot be achieved with all known protocols using local operations and classical communication (LOCC) and linear optics in general, which contradicts the special case discussed in [15], where only point-like sources were considered.
Similarly, in [16], another (LOCC) protocol utilizing N-copy SPE was proposed, and a Fisher information ratio $\mathsf{F}=\frac{N}{N+1}$ was computed. However, upon closer inspection of the protocol we note that it is not implementable with linear optics and requires non-linear operations. We refer the reader to the appendix 2 for a detailed discussion on this protocol.
IV.3 KLM-like scheme
In the following, we introduce several novel quantum telescopy schemes inspired by the Knill-Laflamme-Milburn scheme studied in linear optical quantum computation [22]. A detailed discussion of these protocols is provided in the appendix 3.
KLM scheme[22]:
$|\psi\rangle^{\text{KLM}}_{a}=\frac{1}{\sqrt{N+1}}\sum\limits_{n=0}^{N}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$
The state above has been utilized in the linear optical quantum computation protocol proposed in [22], with which we have from Proposition 1 that:
$\mathbb{h}[|\psi\rangle_{a}^{\text{KLM}}]=\frac{N}{N+1},$ (20)
which scales the same as the $N$-copy SPE, however, this bound is achievable using LOCC and linear optics as we discussed in section V.
Modified-KLM1:
$|\psi\rangle^{\text{Itri}}_{a}=\sum\limits_{n=0}^{N}f^{\text{Itri}}_{n}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$
where $|f^{\text{Itri}}_{n}|^{2}\propto\frac{N}{2}-|\frac{N}{2}-n|$, i.e., the intensity of different terms shaped as a triangle. The QFI ratio is given as:
$\mathbb{h}[|\psi\rangle^{\text{Itri}}_{a}]\stackrel{{\scriptstyle\text{Large {\hbox{N}}}}}{{\approx}}1-2\frac{\log(N)}{N^{2}},$ (21)
By modifying the amplitude of the ancilla state, we could get a even greater QFI ratio asymptotically using the state that has been discussed in [32]
Modified-KLM2[32]:
$|\psi\rangle^{\text{Atri}}_{a}=\sum\limits_{n=0}^{N}f^{\text{Atri}}_{n}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$
where $|f^{\text{Atri}}_{n}|\propto\frac{N}{2}-|\frac{N}{2}-n|$, i.e., the amplitude of different terms shaped as a triangle. The QFI ratio using the above state is given as:
$\mathbb{h}[|\psi\rangle^{\text{Atri}}_{a}]\stackrel{{\scriptstyle\text{Large {\hbox{N}}}}}{{\approx}}1-12\frac{1}{N^{2}}\quad\quad$ (22)
Optimal-KLM:
$|\psi\rangle^{\text{Sub}}_{a}=\sum\limits_{n=0}^{N}\frac{\sqrt{2}\sin(\frac{n+1}{N+2}\pi)}{\sqrt{N+2}}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$
The Optimal-KLM ancilla state is inspired by the upper bound we studied in Theorem 1 discussed in the appendix C, with which, we could get a QFI ratio that almost saturates the numerical optimal (see Fig. 2)
$\mathbb{h}[|\psi\rangle^{\text{Sub}}_{a}]\stackrel{{\scriptstyle\text{Large {\hbox{N}}}}}{{\approx}}1-\pi^{2}\frac{1}{N^{2}}$ (23)
Different from the $N$-copy SPE scheme, We note that all the ‘KLM’-schemes mentioned above can be implemented with linear optics[22] as we will explain in the next section. The implementation is a modified version of the teleportation trick used in linear optical quantum computation, where the state being teleported is one part of a bipartite source state. The encoding is a single-rail encoding[33] with the qubit $|0\rangle$ and $|1\rangle$ representing the vacuum state and the single-photon state, respectively. Before delving into this, we first explain several other protocols, including continuous-variable (CV) teleportation and two-photon entangled states, within the SSR framework we propose.
IV.4 CV teleportation
The continuous-variable-teleportation-based protocol uses the two-mode squeezed vacuum (TMSV) state
$|\psi\rangle^{\text{CV}}_{a}=\frac{1}{\cosh(r)}\sum\limits_{n=0}^{N}\tanh^{n}{r}|n\rangle_{A}|n\rangle_{B}$
to achieve quantum-enhanced telescopy, where the squeezing parameter $r$ is related to the total average photon number $\langle N\rangle=2\sinh^{2}(r)$.
At first glance, from Proposition 1, one can show that $\mathbb{h}[|\psi\rangle^{\text{CV}}_{a}]=0$, which contradicts the protocol introduced in [17, 18] but aligns with the argument made in [34]. As we will explain below, this apparent discrepancy arises from a common misunderstanding in discussions of CV quantum teleportation [35, 36], necessitating a closer examination of the actual resources employed in these protocols.
The apparent contradiction can be resolved by recognizing that, in standard continuous-variable quantum teleportation, an additional pair of correlated coherent states $|\alpha\rangle_{A}|\alpha\rangle_{B}$ , must always be shared. These states ensure that a homodyne measurement can be performed on one side, and a displacement operator can be applied on the other side [17]. This extra resource serves as a phase reference for the CV-teleportation protocol and has also been described as the resource that activates entanglement in a two-mode squeezed state [37]. Consequently, the two-mode squeezed vacuum state and the shared coherent state should be grouped together to define the following state as the actual ancilla used in the CV-teleportation protocol:
$|\psi\rangle^{\text{CV}}_{a}=\frac{1}{\cosh(r)}\sum\limits_{n=0}^{N}\tanh^{n}{r}|n\rangle_{A}|n\rangle_{B}|\alpha\rangle_{A}|\alpha\rangle_{B},$ (24)
As one can numerically verify, the state above has a non-vanishing Quantum Fisher information according to Proposition 1. To avoid the messy calculation with $|\alpha|\rightarrow\infty$ in a CV-teleportation protocol, here we use the Fisher information derived previously [17] to lower bound the Quantum Fisher information ratio of the CV-teleportation protocol, which yields:
$\mathbb{h}[|\psi\rangle^{\text{CV}}_{a}]\geq\mathbb{f}[|\psi\rangle^{\text{CV}}_{a}]\stackrel{{\scriptstyle\text{Large {\hbox{\langle N\rangle}}}}}{{\approx}}1-\frac{1}{\epsilon\langle N\rangle},$ (25)
In the preceding subsections, we discussed several distinct classes of ancilla states where the Quantum Fisher Information (QFI) ratio can asymptotically reach unity. These ancilla states share a crucial characteristic: they simultaneously function as both a nonlocal resource and a phase reference resource [19]. In the following, we present two commonly confused examples of ancilla states used in quantum telescopy and explain why they are unsuitable for achieving quantum-enhanced phase estimation. Specifically, these states serve either solely as a nonlocal (NL) resource or solely as a phase reference (PR) resource.
IV.5 Two-particle entanglement
Compared to single-particle entanglement, two-particle entanglement has been used more frequently in quantum information science, which is represented by the two-particle entangled state[33]
$|\psi\rangle^{\text{TPE}}_{a}=\frac{1}{\sqrt{2}}(|10\rangle_{A}|01\rangle_{B}+|01\rangle_{A}|10\rangle_{B}).$
[…]
Nonetheless, there is a straightforward method to activate the phase reference property of the two-particle ancilla state: by discarding the second mode on both telescopes, the state effectively reduces to the single-particle entangled state used in GJC’s scheme, $|\psi\rangle^{\text{TPE}}_{a}\Rightarrow|\psi\rangle^{\text{SPE}}_{a}$.
IV.6 Correlated coherent state
Another interesting ancilla state to be considered within our framework is the correlated coherent state $|\psi\rangle^{\text{Coh}}_{a}=|\alpha\rangle_{A}|\alpha\rangle_{B}$, which represents states generated from synchronized local oscillators that have been employed to perform the interferometric homodyne measurement in the classical astronomical imaging [7]. It is not difficult to see that for a coherent state, the QFI ratio $\mathbb{h}[|\psi\rangle^{\text{Coh}}_{a}$] is not vanishing. However, this doesn’t contradict the fact that coherent state (homodyne measurement) can only achieve a Fisher information scale $\epsilon^{2}$[11] for a weak thermal source. This is because the QFI calculation serves merely as an upper bound for the achievable Fisher information (FI).
This observation also highlights the fact that the correlated coherent state is not a nonlocal (NL) resource. Furthermore, it serves as an example illustrating that a “local scheme” cannot saturate the Quantum Fisher Information (QFI) in a nonlocal scenario, as demonstrated in other discussions [38]. Similar to the discussion in [11] and as we explain in detail in Appendix 7, we show that, in general, if $\rho_{A}^{T_{A}}>0$ (i.e., the state has a positive partial transpose), the Fisher Information (FI) ratio will always approach zero. This indicates that, in this particular example, it is the lack of nonlocality (entanglement) in the state—rather than its phase reference property—that limits its performance in the quantum telescopy.
Figure 3: Schematic of near-deterministic teleportation: (1) Quantum Fourier transformation $\mathcal{F}_{N+1}$ is implemented at telescope A; (2) Measurement outcome $n$: number of photons detected, and $\vec{d}$: arrangement of detection is sent to telescope B for phase correction; (3) A scanning interferometric measurement is performed at telescope B.
V towards repeater-based linear-optical implementation
We give a short description of a conceivable linear optical implementation of our protocol to illustrate our methodology, with a more detailed derivation in the appendix E. This protocol is an idealisation, as it assumes the loss-free distribution of the ancilla state, such as provided by a repeater-network. The implementation is illustrated in Fig. 3, and operates as follows:
  • The following entangled ancilla state is shared between telescopes A and B[39]:
    $|\psi\rangle_{a}=\sum_{n=1}^{N}f_{n,N-n}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$ (26)
  • A quantum Fourier transformation[40] is performed on $N+1$-modes at telescope A, followed by photon-number resolving measurement at each output port indexed from $[0,\cdots N]$.
  • The measurement outcome at telescope A is characterized by two parameters $(n,\vec{d})$ with $n$ being the total number of photons detected ($n\leq N+1$ since only $N+1$ photon are used in total) and $\vec{d}$ is a $n$-vector includes information about the port number in which each photon is detected, e.g., with the first $3$ photons detected in port 1, the forth photon photon detected in port 2, and no photon in the rest of ports, we have $n=4$ and $d=[1,1,1,2]$.
  • The measurement outcome $(n,\vec{d})$ is sent from telescope A to telescope B. telescope B adds a conditional phase $e^{i\phi_{\vec{d}}}=\omega^{\sum{d}_{i}}$ to its the $n$-th mode of the ancilla at the $n$-th port state when $n\leq N$ (if $n=N+1$, the teleportation protocol fails), where $\omega=\exp[\frac{2\pi i}{N+1}]$.
  • At telescope B, a direct interference measurement is performed with an unbalanced beam splitter with two inputs being the source mode and the $n$-th mode of the telescope B’s part of the ancilla state.
Intuitively, this approach teleports one part of the bipartite source state from telescope A to telescope B. Since neither locality nor superselection rule is a restriction in this protocol, the optimal quantum telescope can be achieved asymptotically with linear optics. Note that the LOCC-based protocol may require a free space delay line at telescope B to compensate for the time required for classical information transmission. However, such a delay line could be made more resilient against noise in a lab compared to the optical channel interconnecting the two telescopes. It could be based, for example, on a broadband all-optical delay-line quantum memory[41] or vacuum delay line [42].
With the advent of quantum repeaters, the lossless distribution of various types of entangled states across nodes in a quantum network has become a promising possibility[43, li2024generalized, Miguel2023]. Similar to the original quantum telescopy protocol[10], a quantum repeater network addresses two key challenges: phase noise, caused by path length variations in interferometers, can be mitigated through active stabilization or entanglement distillation, while photon loss in optical channels can be managed using the repeater protocols. These advancements lay the foundation for the robust implementation of entangled-state distribution in future quantum networks and enable the reliable realization of the quantum telescopy protocols proposed in this letter.
Note that, the implementation above only works for ancilla states in Eq. 26, while the $N$-copy SPE state is not of this particular form with the same mode configurations, we discussed this issue in detail with comments on two related works[14, 15] in the Appendix 2. We want to emphasize that allthough it was argued in both papers that they have proposed linear optical implementation and achieved a Fisher information that seemingly saturates the Quantum Fisher information bound derived in the previous section, we point out some overlooked facts in both papers, see Appendix 2. Our numerical analysis indicates that the scaling of Fisher information ratio is rather $\mathbb{f}\stackrel{{\scriptstyle\text{Large $N$}}}{{\approx}}1-\frac{2}{N}$ with linear optics. And this happens not because of the LOCC constraints on the measurement, but because linear optics can only become universal with extra resources[44, 45].
VI Towards a quantum-enhanced telescopy in noisy repeater-less networks
While future quantum networks promise loss-free entanglement distribution using quantum repeaters [43] in the long-term, it is of interest to study noisy repeater-less quantum telescopy protocols, where the transmission loss for the ground-based entangled photons must be considered. Note that in this scenario the original Gottesman’s scheme will show no benefit over direct transmission, because the ancilla state consists of only one single photon and would suffer the same photon losses as the astronomical photon would. However, more complex entangled-assisted protocols that use many photons could become robust against noise, as shown in [15]. Although this work demonstrated that using an N-SPE state improved phase estimation for an ideal point source (characterized by a mutual coherence function with $|g|=1$ in Eq. 3), as we have discussed in Appendix 2, their result can not be trivially generalized to an extended source (which has mutual coherent function with $|g|<1$ in Eq. 3), which indicates that whenever the source photon is not ideal, the quantum enhancement achieved with $N$-copy SPE state will be degraded.
Looking ahead, it remains unclear which ancilla states are optimal for a noisy network and which measurement protocol could be the more efficient and more compatible with repeaterless linear-optical network. Using our formalism, however, we can determine a fundamental bound for the performance of a quantum telescopy protocol in terms of the average photon number $\langle N\rangle$ in our Corollary 1. When $\langle N\rangle\ll 1$, which is typical for a photon source transmitted after noisy optical channels [46], this bound can be approximated as:
$\mathbb{h}[|\psi\rangle^{\text{CV}}_{a}]\stackrel{{\scriptstyle\text{{\hbox{\langle N\rangle\ll 1}}}}}{{\approx}}\frac{\langle N\rangle}{2},$ (27)
Thus, the primary goal in a realistic quantum telescopy implementation should be to improve the number of photons that survive a lossy channel. Beyond that, the QFI-bound cannot always be saturated with LOCC, especially when the transmitted states experience loss. Therefore, an important next step is to investigate practical quantum telescopy protocols that consider the presence of loss, as expected in near-term and medium-term quantum networks.
We would like to clarify that the primary focus of this letter is to provide a systematic comparison of entanglement-assisted quantum telescopy protocols in a quantum network that operates without repeaters and quantum memories. In this context, each temporal mode emitted by the source is treated as an independent system that cannot be stored, manipulated, or jointly measured. Consequently, the ancilla photons prepared in the laboratory must be supplied individually for each temporal mode. This approach contrasts with memory-based protocols [12, 13], which are expected to become more relevant in the future when fully functional quantum memories are available.
VII Conclusion
We revisited entanglement-assisted telescopy protocols by taking the superselection rules (SSR) into account, which in contrast to the locality constraint, has been broadly overlooked up to now. We argue that the systematic analysis of the SSR constraint is crucial for understanding and characterizing different quantum telescopy protocols and could help constructing new ancilla-state-assisted protocols. Consequently, any entangled-acilla state must be a phase reference and nonlocal resource at the same time to over come these limitations. We indentify several new entanglement-assisted protocols that asymptotically achieve optimal quantum telescopy, which are compatible with linear optical implementations.
Moreover, we closely connect our new ancilla-assisted quantum telescopy protocol to the linear optical quantum computation (LOQC) protocols proposed decades ago. In recent years, numerous new graph-state-based LOQC protocols have emerged, paving the way for multiple generations of such protocols to be implemented in quantum telescopy protocols, particularly once a global quantum network is employed[47].
Several open questions remain, including how to better understand and design practical quantum-enhanced telescopy schemes for near-term quantum networks under imperfections such as ground-photon transmission losses and mode indistinguishability. A recent analysis for the N-SPE protocol[modak2024] assuming an idealized point source, showed a strong degradation of the phase estimation with imperfections, and the performance of our scheme under the same imperfections will be of great interest. Additionally, generalizing our approach to multi-telescope scenarios, again constrained by photon transmission losses and other imperfections, is another important open problem that will be the focus of future work.
Acknowledgments – This work was supported by NSERC Alliance (Hyperspace consortium), NSERC Discovery, and the Government of Canada (ISED) funding to the Institute for Quantum Computing. The authors thank Virginia Lorenz, Yunkai Wang, Eric Chitambar, Raymond Laflamme. Pieter Kok and Robert Spekkens for helpful discussions.
Appendix
[…]
1
Appendix A Thermal source state and optimal quantum Fisher information
In quantum optics, the astronomical object viewed by two telescopes can be modeled as a two-mode quasi-monochromatic thermal source[11, 23]:
$\rho_{s}=\iint d^{2}\alpha d^{2}\beta\frac{1}{\pi^{2}\det\Gamma}\exp[-(\alpha^{*},\beta^{*})\Gamma^{-1}\begin{pmatrix}\alpha\\ \beta\end{pmatrix}]|\alpha,\beta\rangle\langle\alpha,\beta|$ (1)
where $\Gamma=\frac{\epsilon}{2}\begin{pmatrix}1&g\\ g^{*}&1\end{pmatrix}$ and $g=|g|e^{i\theta}$ is the complex visibility, also known as, mutual coherent function or first-order spatial coherence. $\epsilon$, which quantifies the strength of the source, is typically small for astronomical sources at optical wavelength[5]. In the limit of $\epsilon\ll 1$, we can write the density operator of the state of the optical field on the image plane as
$\rho_{s}=(1-\epsilon)\rho_{s}^{(0)}+\epsilon\rho_{s}^{(1)}+O(\epsilon^{2})\rho_{s}^{(>1)},$ (2)
Where the leading order term encoded with the complexity visibility (coherence function) $g=|g|e^{i\theta}$ is the single-photon term $\rho_{s}^{(1)}$, which can be written in the telescope Fock basis $\{|0\rangle_{A}|1\rangle_{B},|1\rangle_{A}|0\rangle_{B}\}$ as:
$\rho_{s}^{(1)}=\frac{1}{2}[|0_{A}1_{B}\rangle\langle 0_{A}1_{B}|+|1_{A}0_{B}\rangle\langle 1_{A}0_{B}|+(g|1_{A}0_{B}\rangle\langle 0_{A}1_{B}|+g^{*}|0_{A}1_{B}\rangle\langle 1_{A}0_{B}|)]$ (3)
The van Cittert-Zernike theorem [2] states that the visibility (as a function of baseline) is the Fourier transform of the source distribution. Thus, the primary goal in astronomical interferometric imaging is to make a good estimate of complexity visibility for all baselines.
From the quantum estimation theory, the amount of information that is contained in the state in Eq. 2 can be quantified by the corresponding quantum Fisher information[24]. Focus on the single photon term $\rho_{1}$, the symmetric logarithmic derivative of it is a solution of the differential equation and is given as:
$\frac{\hat{L}_{\mu}\rho_{s}+\rho_{s}\hat{L}_{\mu}}{2}=\frac{\partial\rho_{s}}{\partial\mu}\quad\quad\mu\in\{|g|,\theta\}\quad\quad\Longrightarrow\quad\quad\hat{L}_{|g|}=\frac{1}{1-|g|^{2}}\begin{pmatrix}-g&e^{i\theta}\\ e^{-i\theta}&-g\end{pmatrix}\quad\quad\hat{L}_{\theta}=|g|\begin{pmatrix}0&ie^{i\theta}\\ -ie^{-i\theta}&0\end{pmatrix}$ (4)
which yields the optimal QFI of estimating $|g|$ and $\theta$:
$\hat{H}_{|g|}[\rho_{s}]=\textrm{Tr}[\rho_{s}\hat{L}^{2}_{|g|}]=\frac{\epsilon}{1-|g|^{2}}+O(\epsilon^{2})~{}~{}~{}~{}~{}~{}\quad\quad\quad\hat{H}_{\theta}[\rho_{s}]=\textrm{Tr}[\rho_{s}\hat{L}^{2}_{\theta}]=|g|^{2}\epsilon+O(\epsilon^{2})$ (5)
However, the quantum Fisher information is not always achievable with some, especially if the system is under-constrained and not all possible measurement can be implemented. The typical constraint that has been recognized in the past is the locality constraint, and as has been discussed by Tsang in[11], all local measurements performed using local operations and classical communication (LOCC), cannot achieve the optimal QFI. In this letter, we will focus on a different constraint that exists in any interferometric setup, known as the superselection rule (SSR). Specifically, there are no shared phase references, the allowed state preparation and measurement are constrained by certain symmetry due to superselection rule as we will explain later. By studying this type of constraint, we can answer different questions that taking locality constraints alone cannot tackle. For example, we will be able to compare the power of different entanglement-assisted protocols within this framework and present several novel schemes that asymptotically achieve the optimal quantum protocol, i.e., obtain optimal quantum Fisher information asymptotically in the most resource-efficiently way (consuming less resources).
Appendix B Superselection rule, reference frame, and quantum estimation under SSR
In this section, we will discuss in detail the photon-number superselection rule encountered in the interferometric imaging system and its impact on the quantum estimation problem. We will start by giving a short textbook explanation of the reference frame and superselection rule. In this letter, we will focus only on the phase reference (PR) frame and photon-number $U(1)$ superselection rule, and we refer readers to the review paper by Bartlett, Rudolph, and Spekkens[26] for more discussions:
[…]
In a quantum optics experiment, the quantum states of optical modes are always defined relative to some phase reference[19]. Consider a state $|\psi\rangle$ defined relative to Alice phase reference frame as $|\psi\rangle=\sum_{n=0}c_{n}|n\rangle$ in the photon number basis. However, the state will be described differently in different phase references, say in Bob’s phase references, by a unitary transformation $U(\phi)=e^{i\phi\hat{N}}$, where $\phi$ is the angle between Bob’s phase references and Alice’s phase references, and $\hat{N}$ is the number operator. More specifically, the state viewed by Bob becomes:
$|\psi\rangle=\sum_{n=0}c_{n}|n\rangle\quad\quad\Longrightarrow\quad\quad U(\psi)|\phi\rangle=\sum_{n=0}e^{in\phi}c_{n}|n\rangle$ (6)
Now consider the situation where the phase differences $\phi$ are completely unknown by Bob or Alice, the lack of phase information restricts the kind of state prepared by Alice that can be recognized by Bob, more religiously, any state $|\psi\rangle$ prepared by Alice, becomes averaged out due to the lack of phase $\phi$ as
$|\psi\rangle=\sum_{n=0}c_{n}|n\rangle\quad\Longrightarrow\quad\mathcal{U}(\rho):=\int_{0}^{2\pi}U(\psi)|\phi\rangle\langle\phi|U(\phi)^{\dagger}=\sum_{n,n^{\prime}=0}\int_{0}^{2\pi}\frac{d\phi}{2\pi}e^{i(n-n^{\prime})\phi}c_{n}c^{*}_{n}|n^{\prime}\rangle\langle n^{\prime}|=\sum_{n=0}|c_{n}|^{2}|n\rangle\langle n|$ (7)
Therefore, the lack of phase references is mathematically equivalent to the so-called photon-number selection rules, where state preparation, manipulation, and discrimination have coherence between photon-number basis is forbidden, this superselection rule is also known as $U(1)$-superselection rule, in this letter, we will use photon-number SSR, $U(1)$-SSR and lack of phase references interchangeably.
In a theory where $U(1)$-SSR is imposed, it is useful to decompose the Hilbert space $H$ as the direct sum of number-eigen subspace:
$H=\bigoplus_{n=0}^{N}H_{n}\quad\quad\quad\text{with }H_{n}=P_{n}HP_{n},$ (8)
with $P_{n}$ being the projector onto the $n$-photon subspace, i.e., $P_{n}=|n\rangle\langle n|$. Now, consider a bipartite system of Alice and Bob, where Alice and Bob describe their quantum mechanics in their Hilbert space $H_{A}$ and $H_{B}$, If third-party want to describe the bipartite quantum systems, the $U(1)$-SSR rule force the Hilbert space $H=H_{A}\otimes H_{B}$ to be decomposed as[27, 28, 29]
$H=\bigoplus_{N=0}^{\infty}H_{N}=\bigoplus_{N=0}^{\infty}\bigoplus_{n=1}^{N}H^{A}_{n}\otimes H^{B}_{N-n}$ (9)
Effectively, SSR imposed on certain constrain a bipartite system, which gives:
Lemma 1.
A bipartite state $\rho^{AB}$ under photon-number superselection rules will be represented as[29]:
$\rho^{AB}=\mathcal{E}(\rho^{AB})=\sum_{n,m=0}^{\infty}(P^{A}_{n}\otimes P^{B}_{m})\rho^{AB}(P^{A}_{n}\otimes P^{B}_{m})$ (10)
where $P^{A}_{n}=|n\rangle\langle n|$ represent projector onto $H^{A}_{n}$, and similarly for $P^{B}_{n}$
Proof.
A superselection rule imposed on the bipartite quantum state $\rho^{AB}$ implies a lack of shared phase references on the bipartite system. Since the phase references on systems A and B are uncorrelated, the mixing over all different phases is described as:
$\mathcal{E}(\rho^{AB})$ $=\iint_{0}^{2\pi}\frac{d\phi_{1}}{2\pi}\frac{d\phi_{2}}{2\pi}\mathcal{U}^{A}(\phi_{1})\otimes\mathcal{U}^{B}(\phi_{2})[\rho^{AB}]$
$=\sum_{n,m,n^{\prime},m^{\prime}=0}\iint_{0}^{2\pi}\frac{d\phi_{1}}{2\pi}\frac{d\phi_{2}}{2\pi}e^{i(n-n^{\prime})\phi_{1}+i(m-m^{\prime})\phi_{2}}(P^{A}_{n}\otimes P^{B}_{m})\rho^{AB}(P^{A}_{n^{\prime}}\otimes P^{B}_{m^{\prime}})$
$=\sum_{n,m=0}^{\infty}(P^{A}_{n}\otimes P^{B}_{m})\rho^{AB}(P^{A}_{n}\otimes P^{B}_{m})$ (11)
where $\mathcal{U}(\phi)[\rho]=e^{in\phi}=\sum_{n}e^{in\phi}P_{n}\rho P_{n}e^{-in\phi}$ is also known as $U(1)$-twirling operation[48, 26]. ∎
Next, we look at the consequence of the superselection rule in the estimation problem that we are interested in. A direct consequence of lemma 10 implies that the source under consideration in Eq. 2 will contain no information on the complexity visibility $g$ for its zero-order and first-order terms. i.e.,
$\mathcal{E}(\rho_{s})=(1-\epsilon)\mathcal{E}(\rho_{s}^{(0)})+\epsilon\mathcal{E}(\rho_{s}^{(1)})+O(\epsilon^{2})\mathcal{E}(\rho_{s}^{(>1)})=(1-\epsilon)\mathbb{I}^{(0)}+\epsilon\frac{\mathbb{I}^{(1)}}{2}+O(\epsilon^{2})\mathcal{E}(\rho_{s}^{(>1)}),$
where $\mathbb{I}^{n}$ is the identity operator in the $n$-party subspace of $H_{n}$. Therefore, in the presence of SSR the achievable QFI always scales at most $\epsilon^{2}$, which explains why an intensity interferometer is not preferred in very-long baseline interferometry[30, 5].
However, it is always possible to use a preshared bipartite state to build shared phase references and lift the restriction of SSR[49], this idea is also studied and known as activating entanglement under SSR[50]. Simply put, by introducing an additional ancillary system $\rho_{a}$, the composite system of $\rho^{AB}=\rho^{AB}_{a}\otimes\rho^{AB}_{s}$ could contain more information about the complexity visibility $g$ that is not erased by the superselection rules.
More rigorously, when an ancilla state is included, a composite system takes the following form under the superselection rule:
$\mathcal{E}(\rho^{AB})=\sum_{n,m=0}^{\infty}(P^{A}_{n}\otimes P^{B}_{m})\rho_{s}^{AB}\otimes\rho_{a}^{AB}(P^{A}_{n}\otimes P^{B}_{m})$ (12)
Since $\rho_{a}^{AB}$ have no dependence on $g$, the additivity and Monotonicity of QFI guarantee that:
Lemma 2.
The Quantum Fisher information of the composite system in the presence of SSR is upper bounded by the optimal one
$\mathbb{H}_{|g|}[\mathcal{E}(\rho^{AB})]\leq\mathbb{H}_{|g|}[\rho_{s}^{AB}]\quad\quad\mathbb{H}_{\theta}[\mathcal{E}(\rho^{AB})]\leq\mathbb{H}_{\theta}[\rho_{s}^{AB}]$ (13)
Proof.
From the additivity and Monotonicity of quantum Fisher information, we have
$\mathbb{H}_{\mu}[\rho_{s}^{AB}]+\mathbb{H}_{\mu}[\rho_{a}^{AB}]\geq\mathbb{H}_{\mu}[\rho_{s}^{AB}\otimes\rho_{a}^{AB}]\geq\mathbb{H}_{\mu}[\mathcal{E}(\rho^{AB})]\quad\quad\mu\in\{|g|,\theta\}$ (14)
Since $\mathbb{H}_{\mu}[\rho_{a}^{AB}]$=0, we have the lemma proved. ∎
This motivates us to define
$\mathbb{h}[\rho^{AB}_{a}]=\min\{\frac{\mathbb{H}_{|g|}[\mathcal{E}(\rho^{AB})]}{\mathbb{H}_{|g|}[\rho_{s}^{AB}]},\frac{\mathbb{H}_{\theta}[\mathcal{E}(\rho^{AB})]}{\mathbb{H}_{\theta}[\rho_{s}^{AB}]}\}$ (15)
which takes value in $[0,1]$ from lemma 13. In the following, we will refer to this quantity as the QFI ratio, and work out finding the optimal entanglement ancilla state such that the QFI ratio can asymptotically reach one, thus achieving optimal quantum telescopy.
Appendix C Upper Bound of Quantum Fisher Information Ratio.
Proposition 1.
Given an ancilla state $\rho_{a}=|\psi\rangle\langle\psi|_{a}$ of the form in Eq. 8, the quantum fisher information in estimating $|g|$ and $\theta$ of source state $\rho_{s}$ in the presence of superselection rule is given as:
$\mathbb{H}_{|g|}[\mathcal{E}(\rho^{AB})]=\frac{\epsilon}{1-|g|^{2}}\sum_{n,m=1}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}$
$\mathbb{H}_{\theta}[\mathcal{E}(\rho^{AB})]=\epsilon|g|^{2}\sum_{n,m=1}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}$
$\mathbb{h}[\rho^{AB}_{a}]=\sum_{n,m=1}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}$ (16)
Here we neglect the higher term $O(\epsilon^{2})$ for simplicity. So in the following $\mathbb{h}=0$ actually means $\mathbb{h}=O(\epsilon^{2})$.
Proof.
By the convexity of quantum Fisher information, it suffices to consider pure ancilla state $\rho_{a}=|\psi\rangle\langle\psi|_{a}$ when optimizing all ancill-assisted protocols. The shared ancilla state can be written in the most general form:
$|\psi\rangle_{a}=\sum_{n,m=0}f_{n,m}|n_{A},m_{B}\rangle,$ (17)
with $\sum_{n,m=0}|f_{n,m}|^{2}=1$. We abuse $|n_{A},m_{B}\rangle$ to represent a state without specifying its mode configuration but any state with $n$ photons at telescope $A$ and $m$ photons at telescope $B$, for example, it could represent
  • $|n_{A},m_{B}\rangle=|n\rangle_{A}|m\rangle_{B}$: A product state of n photon in one mode
  • $|n_{A},m_{B}\rangle=|1\rangle_{A}^{\otimes n}|0\rangle_{A}^{\otimes m}|0\rangle_{B}^{\otimes n}|1\rangle_{B}^{\otimes m}$: A product state $(n+m)$ modes at both $A$ and $B$, where there exist 1 photon per mode at the first $n$ mode at A, and 1 photon per mode at the last $m$ mode at B
  • $|n_{A},m_{B}\rangle=\sum_{\pi}\sqrt{p(\pi)}\pi(|1\rangle_{A}^{\otimes n}|0\rangle_{A}^{\otimes m}|0\rangle_{B}^{\otimes n}|1\rangle_{B}^{\otimes m})$, where $\pi$ is different permutation on the modes and the state is no longer a product state but an entangled state.
Since the specific mode configuration $|n_{A},m_{B}\rangle$ of the ancilla state does not affect the quantum Fisher information of the composite system, we will not specific the mode configuration until we are discussing the specific protocols.
By neglecting the higher order term from $\rho_{s}$, we have:
$\mathcal{E}(\rho^{AB})=\sum_{n,m}(P^{A}_{n}\otimes P^{B}_{m})\rho_{s}^{AB}\otimes\rho_{a}^{AB}(P^{A}_{n}\otimes P^{B}_{m})=\bigoplus_{n,m}\sigma^{AB}_{n,m}$ (18)
where $\sigma^{AB}_{n,m}$ are normalized states with $n$-photon at telescope A, and $m$-photon at telescope B, which takes the form of:
$\sigma^{AB}_{n,m}$ $=(1-\epsilon)|f_{n,m}|^{2}|0,n_{A}\rangle\langle 0,n_{A}|\otimes|0,m_{B}\rangle\langle 0,m_{B}|+\frac{\epsilon}{2}\left([f_{n,m-1}|^{2}|0,n_{A}\rangle\langle 0,n_{A}|\otimes|1,(m-1)_{B}\rangle\langle 1,(m-1)_{B}|\right.$ (19)
$+f_{n-1,m}f^{*}_{n,m-1}g|0,n_{A}\rangle\langle 1,(n-1)_{A}|\otimes|1,(m-1)_{B}\rangle\langle 0,m_{B}|+|f_{n-1,m}|^{2}|1,(n-1)_{A}\rangle\langle 1,(n-1)_{A}|\otimes|0,m_{B}\rangle\langle 0,m_{B}|$
$+f^{*}_{n-1,m}f_{n,m-1}g^{*}|1,(n-1)_{A}\rangle\langle 0,n_{A}|\otimes|0,m_{B}\rangle\langle 1,(m-1)_{B}|\left.\right)+O(\epsilon^{2})$ (20)
Where $|0,n_{A}\rangle$ is state with 00 photon from the ancilla and $n_{A}$ photon from the source. Since $\sigma_{n,m}^{A,B}$ are density operator defined on different orthogonal Hilbert spaces, it is not difficult to see that:
$\mathbb{H}_{\mu}[\mathcal{E}(\rho^{AB})]=\sum_{n,m=0}\mathbb{H}_{\mu}[\sigma_{n,m}^{AB}]\quad\quad\quad\mu\in\{|g|,\theta\}$ (21)
where the QFI for the individual unnormalized state is discussed below in lemma 3 and is given as:
[…]
Lemma 3.
The quantum Fisher information contains in the unnormalized state $\sigma_{nm}^{AB}$ is contributed by the second term in Eq. 20, which can be rewritten as:
$\sigma^{AB\epsilon}_{n,m}=\frac{(|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2})\epsilon}{2}\begin{pmatrix}\frac{|f_{n,m-1}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}&\frac{f_{n-1,m}f^{*}_{n,m-1}g}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}g\\ \frac{f^{*}_{n-1,m}f_{n,m-1}g}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}g^{*}&\frac{|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}\end{pmatrix}$ (23)
in $\{|0,n_{A}\rangle|1,(m-1)_{B}\rangle,|1,(n-1)_{A}\rangle|0,m_{B}\rangle\}$ basis. And the normalized version of it is the matrix term above, donated as $\tilde{\sigma}^{AB\epsilon}_{n,m}$ is the term inside the Bracket. Similar to the calculation in section A, one could show that the symmetric logarithmic derivative of $\tilde{\sigma}^{AB\epsilon}_{n,m}$ in estimating $|g|$ and $\theta$ is given as
$\hat{L}_{|g|}[\tilde{\sigma}^{AB\epsilon}_{n,m}]$ $=\frac{1}{1-|g|^{2}}\frac{2}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}\begin{pmatrix}-|f_{n-1,m}|^{2}g&f_{n-1,m}f^{*}_{n,m-1}e^{i\theta}\\ f^{*}_{n-1,m}f_{n,m-1}e^{-i\theta}&-|f_{n,m-1}|^{2}g\end{pmatrix}$
$\hat{L}_{\theta}[\tilde{\sigma}^{AB\epsilon}_{n,m}]$ $=|g|\frac{2}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}\begin{pmatrix}0&if_{n-1,m}f^{*}_{n,m-1}e^{i\theta}\\ -if^{*}_{n-1,m}f_{n,m-1}e^{-i\theta}&0\end{pmatrix}$
Therefore, we have:
$\mathbb{H}_{|g|}[\sigma_{n,m}^{AB}]$ $=\frac{(|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2})\epsilon}{2}\mathbb{H}_{|g|}[\tilde{\sigma}^{AB\epsilon}_{n,m}]=\frac{\epsilon}{1-|g|^{2}}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}$
$\mathbb{H}_{\theta}[\sigma_{n,m}^{AB}]$ $=\frac{(|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2})\epsilon}{2}\mathbb{H}_{\theta}[\tilde{\sigma}^{AB\epsilon}_{n,m}]=\epsilon|g|^{2}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}$
Let known discuss shared ancilla state $|\psi\rangle_{a}$ restricted with bounded energy, a natural choice to implement such restriction on the ancilla state is to limit the maximum photon number to a fixed value $N$, with which we have the following theorem.
Theorem 1.
For ancilla state $|\psi\rangle_{a}$ using at most $N$-photon in total, we have:
$\mathbb{h}[\rho^{AB}_{a}]\leq\cos(\frac{\pi}{N+2})\stackrel{{\scriptstyle\text{Large $N$}}}{{\approx}}(1-\frac{\pi^{2}}{2N^{2}})$ (24)
Proof.
Assuming, the ancilla state contributes $N$ photon at most, we should only consider unnormalized state $\sigma_{n,m}^{AB}$ with $n+m\leq N+1$ (with at most $N$ photon from the ancilla state and $1$ photon from the source)
It is convenient to simplify the notation by firstly replacing the summation:
$\mathbb{h}[\rho^{AB}_{a}]=\sum_{n,m=1}^{n+m\leq N+1}\frac{2|f_{n,m-1}|^{2}|f_{n-1,m}|^{2}}{|f_{n,m-1}|^{2}+|f_{n-1,m}|^{2}}\Longrightarrow\mathbb{h}[\rho^{AB}_{a}]=\sum_{k=1}^{N}\sum_{n=1}^{k}\frac{2|f_{n,k-n}|^{2}|f_{n-1,k-n+1}|^{2}}{|f_{n,k-n}|^{2}+|f_{n-1,k-n+1}|^{2}}$ (25)
Looking at each term in the summation on the right hand side, we could define the two normalized probabilities for every $k$ as:
$p^{k}_{n}:=\frac{|f_{n,k-n}|^{2}}{\sum_{n=0}^{k}|f_{n,k-n}|^{2}}\quad\quad\quad\quad q_{k}=\sum_{n=0}^{k}|f_{n,k-n}|^{2}$
where $\sum_{k}q_{k}$ and $\sum_{n}p^{k}_{n}=1$ are normalized representing the probability of having $k$ total photon from the ancilla state $\rho^{AB}_{a}$ and conditional on that, having $n$ photon at telescope A. With this more compact notation, we could show that each term in $\mathbb{h}[\rho^{AB}_{a}]$:
$h_{k}:=q_{k}\sum_{n=1}^{k}\frac{2p_{n}^{k}p_{n-1}^{k}}{p_{n}^{k}+p_{n-1}^{k}}\leq q_{k}\sum_{n=1}^{k}\sqrt{p_{n}^{k}p_{n-1}^{k}}$ (26)
Now, we can give a non-tight upper bound $h_{k}$ by upper bounding $\sum_{n=1}^{k}\sqrt{p_{n}^{k}p_{n-1}^{k}}$ from the lemma below, which shows
$h_{k}\leq q_{k}\cos(\frac{\pi}{k+2})\quad\quad\Longrightarrow\quad\quad\mathbb{h}[\rho^{AB}_{a}]=\sum_{k=1}^{N}h_{k}\leq\sum_{k=1}^{N}q_{k}\cos(\frac{\pi}{k+2})\leq\cos(\frac{\pi}{N+2})$
with the second inequality saturated when $q_{k}=\delta_{k,N}$, which implies:
$f_{n,m}=f_{n,m}\delta_{n,N-m}=:f_{n}\delta_{n,N-m}$
Therefore, in the following discussion, we focus on states of the following form specifically:
$|\psi\rangle_{a}=\sum_{n=0}^{N}f_{n}|n_{A},(N-n)_{B}\rangle,$ (27)
Lemma 4.
For any distribution $\{x_{i}\}$ with $\sum_{i=0}^{k}x_{i}=1$, we have the inequality:
$\sum_{i=1}^{k}\sqrt{x_{i}x_{i-1}}\leq\cos(\frac{\pi}{k+2}),$ (28)
and the inequality is saturated with $x_{i}={\frac{2}{N+2}}\sin^{2}(\frac{i+1}{k+2}\pi)$
Proof.
Defining a tridiagonal matrix $M$ of the form:
$\begin{pmatrix}0&1&0&\cdots&0\\ 1&0&1&\cdots&0\\ 0&1&0&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&1\\ 0&0&\cdots&1&0\end{pmatrix}$ (29)
one could easily see that $2\sum_{i=1}^{k}\sqrt{x_{i}x_{i-1}}$ corresponds precisely to the expectation value of a $k+1$ by $k+1$ matrix $M_{k+1}$ and normalized vector $\vec{v}=(\sqrt{x_{0}},\cdots,\sqrt{x_{k}})^{T}$. Therefore, to find upper bound on $2\sum_{i=1}^{k}\sqrt{x_{i}x_{i-1}}$ for all distribution $\{x_{i}\}$, it is equivalent to compute the maximal eigenvalue of $M_{k+1}$, which has been discussed in previous work [51, 52]. The determinant of $M_{k+1}$ can be expanded as:
$\det(M_{k+1}-\lambda I_{k+1})=-\lambda\cdot\det(M_{k}-\lambda I_{k})-\det(M_{k-1}-\lambda I_{k-1})$
where $I_{k}$ are $k$ by $k$ identity matrix. This recursive formula $\det(M_{k}-\lambda I_{k})$ can be related to the recurrence relation of Chebyshev polynomials $U_{k+1}(-\lambda/2)$, with zeros at
$\lambda_{i}=2\cos(\frac{i+1}{k+2}\pi)\quad\quad\quad i\in[0,\cdots k]$
thus, the largest eigenvalue of $M_{k+1}=2\cos(\frac{\pi}{k+2})$ with eigenvector $\vec{v}\propto(\sin(\frac{1}{k+2}\pi),\sin(\frac{2}{k+2}\pi),\cdots,\sin(\frac{k+1}{k+2}\pi))^{T}$
Another way of imposing the constraint could be on the average photon number $\langle N\rangle=\sum_{n,m}|f_{n,m}|^{2}(n+m)=\langle N\rangle$, where a similar bound could be given as:
Corollary 2.
For ancilla state $|\psi\rangle_{a}$ average photon number $\langle N\rangle$, we have:
$\mathbb{h}[\rho^{AB}_{a}]\leq\cos(\frac{\pi}{\langle N\rangle+2})\stackrel{{\scriptstyle\text{Large $\langle N\rangle$}}}{{\approx}}(1-\frac{\pi^{2}}{2\langle N\rangle^{2}})$ (30)
Proof.
Note that the average photon number $\langle N\rangle=\sum_{n,m}|f_{n,m}|^{2}(n+m)=\langle N\rangle$ can be rewritten as:
$\sum_{k=0}\sum_{n=0}^{k}|f_{n,k-n}|^{2}k=\langle N\rangle$
Again defining $q_{k}=\sum_{n=0}^{k}|f_{n,k-n}|^{2}$, and we want solve the following optimization problem
$\text{sup}\sum_{k=0}^{\infty}q_{k}\cos(\frac{\pi}{k+2}\pi)$
[…]
Since $\cos(\frac{\pi}{k+2}\pi)$ is a convex function, from Jensen’s inequality ($E(\Phi(x))\leq\Phi(E(X))$ for any random variable $X$ and convex function $\Phi$):
$\sum_{k=0}^{\infty}q_{k}\cos(\frac{\pi}{k+2})\leq\cos(\frac{\pi}{\langle N\rangle+2})$
And the inequality saturate if $\langle N\rangle$ is an integer and $q_{k}=\delta_{k,\langle N\rangle}$
Appendix D Discussion on different entanglement-assisted protocol
1 Single particle entangled states
We begin by discussing the GJC scheme [10] within our theoretical framework. In the GJC scheme, a single-particle entangled state, $|\psi\rangle_{a}=\frac{1}{\sqrt{2}}(|0\rangle_{A}|1\rangle_{B}+|1\rangle_{A}|0\rangle_{B})$, is first used to realize a quantum telescope. Using Proposition 1, it is straightforward to verify that:
$\mathbb{h}[|\psi\rangle_{a}]=\frac{1}{2}$ (31)
These bounds establish a fundamental limit on quantum estimation problems in quantum telescopes using single-particle entangled states, $|\psi\rangle_{a}$, which has already been demonstrated with the correlated interference measurement (a local measurement) in the original quantum telescopy scheme [10, 11]. It was argued in [14] that nonlinear optics can help improve the Fisher information. This appears to contradict our conclusions, as our upper bound based on quantum Fisher information does not exclude nonlinear-optics-based measurements. However, upon closer inspection, we realize that implementing the control gates used in [14] at two telescopes coherently is fundamentally impossible without utilizing additional resources to bypass the superselection rule. More specifically, under a $U(1)$-superselection rule, not only must the allowed states but also the allowed operations be $U(1)$-invariant [19]. Consequently, two control gates cannot be prepared at two telescopes if their phase references are not aligned.
Although the original GJC scheme cannot be improved by employing different operations or measurements, it is still possible to enhance its performance by using different shared entangled states. Next, we provide a natural extension of the GJC scheme by utilizing $N$-copy single-particle entangled states shared between telescopes.
2 N-copy single particle entangled states
We note that the use of N-copy SPE: $|\psi\rangle^{\otimes N}_{a}=\frac{1}{\sqrt{2^{N}}}(|0\rangle_{A}|1\rangle_{B}+|1\rangle_{A}|0\rangle_{B})^{\otimes N}$ has been previously discussed [16, 15], however, their attempt didn’t fully solve the problem as we will comment on later, more importantly, the performance of N-copy SPE has not been carefully studied so far. Following Proposition 1, for N-copy SPE, we have
$|\psi\rangle^{\otimes N}_{a}=\sum_{n=0}^{N}\sqrt{\frac{{N\choose n}}{2^{N}}}|n_{A},(N-n)_{B}\rangle\Rightarrow\mathbb{h}[|\psi\rangle^{\otimes N}_{a}]=\frac{N}{N+1}$ (32)
Note that the state $|n_{A},(N-n)_{B}\rangle$ is not a product state, but instead entangled of the form:
$|n_{A},(N-n)_{B}\rangle=\frac{1}{\sqrt{{N\choose n}}}\sum_{\pi}\pi(|1\rangle_{A}^{\otimes n}|0\rangle_{A}^{\otimes m}\otimes|0\rangle_{B}^{\otimes n}|1\rangle_{B}^{\otimes m})$ (33)
where $\pi$ represents different permutations on the photon modes (There are $N=n+m$ modes in both telescope A and B, so there are in total (${N\choose n}$ such configurations. ).
Asymptotically, the quantum Fisher information goes to 1, as the number of copies of SPE goes to one. In different special cases, this bound has been saturated using different measurement schemes including the estimating of point source in [15] with Linear optics, and the estimation of a general source with nonlinear operation. Here we will briefly comment on these two papers.
In [16], a phase extraction scheme was proposed in Section. IV, where it was argued that their scheme could be done with only linear optical elements. However, we found out that the operation required to transform Eq. 51 to Eq. 52 in their paper can not be implemented with linear optics. More specifically, linear optical operations are not capable of performing quantum Fourier transformations over different $N$ photon states for $N>1$ (i.e., $\hat{a}_{2}\hat{a}_{3}\nRightarrow\frac{1}{\sqrt{3}}(\hat{a}_{2}\hat{a}_{3}+\hat{a}_{1}\hat{a}_{3}+\hat{a}_{1}\hat{a}_{2})$ is not allowed), since the operations should not only be linear but also be described as a unitary operation on individual creation operator.(i.e., $\hat{a}_{3}\Rightarrow\frac{1}{\sqrt{3}}(\hat{a}_{3}+\hat{a}_{1}+\hat{a}_{2})$ is allowed). However, if one is not restricted with linear optics but just LOCC, the phase extraction scheme introduced in [16] is capable of estimating not only $\theta$ but also $|g|$ thus saturate the Quantum Fisher information ratio we report here.
In [15], a protocol is presented that uses only linear optics to estimate the phase $\theta$ of a point source (with $|g|=1$). This protocol achieves a Fisher information ratio for estimating $\mathbb{f}^{N}_{\theta}$ that scales as $1-\frac{1}{N+1}$ (defined as $\mathbb{F}{\theta}/\mathbb{H}{\theta}$), where $N$ is the number of copies of the shared single-particle entangled state (SPE). In this work, we extend their results by explicitly computing the Fisher information for estimating both $|g|$ and $\theta$ using their scheme for $N\in{1,\cdots,5}$.
$\mathbb{F}_{\theta}^{1}=\frac{1}{2}\frac{\sin^{2}\phi}{1-|g|^{2}\cos^{2}\phi}|g|^{2}\epsilon$ $\quad\mathbb{F}_{|g|}^{1}=\frac{1}{2}\frac{\cos^{2}\phi}{1-|g|^{2}\cos^{2}\phi}\epsilon$
$\mathbb{F}_{\theta}^{2}=\frac{3\sin^{2}\phi}{(1-|g|\cos\phi)(5+4|g|\cos\phi)}|g|^{2}\epsilon$ $\quad\mathbb{F}_{|g|}^{2}=\frac{3\cos^{2}\phi}{(1-|g|\cos\phi)(5+4|g|\cos\phi)}\epsilon$
$\mathbb{F}_{\theta}^{3}=\frac{3\sin^{2}\phi(9+7|g|\cos\phi)}{4(1-|g|^{2}\cos^{2}\phi)(10+6|g|\cos\phi)}|g|^{2}\epsilon$ $\quad\mathbb{F}_{|g|}^{3}=\frac{3\cos^{2}\phi(9+7|g|\cos\phi)}{4(1-|g|^{2}\cos^{2}\phi)(10+6|g|\cos\phi)}\epsilon$
$\mathbb{F}_{\theta}^{4}=\frac{10\sin^{2}\phi(16+9|g|\cos\phi)}{(1-|g|\cos\phi)(13+12|g|\cos\phi)(17+8|g|\cos\phi)}|g|^{2}\epsilon$ $\quad\mathbb{F}_{|g|}^{4}=\frac{10\cos^{2}\phi(16+9|g|\cos\phi)}{(1-|g|\cos\phi)(13+12|g|\cos\phi)(17+8|g|\cos\phi)}\epsilon$
$\mathbb{F}_{\theta}^{5}=\frac{5\sin^{2}\phi(79+106|g|\cos\phi+31|g|^{2}\cos^{2}\phi)}{(1-|g|^{2}\cos^{2}\phi)(26+10|g|\cos\phi)(20+16|g|\cos\phi)}|g|^{2}\epsilon$ $\quad\mathbb{F}_{|g|}^{5}=\frac{5\cos^{2}\phi(79+106|g|\cos\phi+31|g|^{2}\cos^{2}\phi)}{(1-|g|^{2}\cos^{2}\phi)(26+10|g|\cos\phi)(20+16|g|\cos\phi)}\epsilon$ (34)
where $\phi=\delta-\theta$ with $\delta$ being the phase delay used in the measurement. In [15], they showed implicitly that $\phi=0$ is the optimal point to get the optimal Fisher information ratio $\mathbb{f}^{N}_{\theta}=1-\frac{1}{N+1}$ for point source with $|g|=1$, agree with Eq. 34 above, however, we could easily check that $\mathbb{f}^{N}_{\theta}=0$ at $\phi=0$ for a source with $|g|<1$ (which is almost always the case), in fact the Fisher information ratio scaling $1-\frac{1}{N+1}$ is an artifact that occurs because $|g|=1,\phi=0$ is a removable singular point. As some concrete examples, one can simply verify that:
$\mathbb{F}_{\theta}^{2}$ $=\frac{3\sin^{2}\phi}{(1-|g|\cos\phi)(5+4|g|\cos\phi)}|g|^{2}\epsilon\stackrel{{\scriptstyle\text{$|g|=1$, $\phi\rightarrow 0$}}}{{\Longrightarrow}}\frac{3\phi^{2}}{(\frac{1}{2}\phi^{2})(5+4)}\epsilon=\frac{2}{3}|g|^{2}\epsilon$ (35)
$\mathbb{F}_{\theta}^{2}$ $=\frac{3\sin^{2}\phi}{(1-|g|\cos\phi)(5+4|g|\cos\phi)}|g|^{2}\epsilon\stackrel{{\scriptstyle\text{$|g|<1$, $\phi\rightarrow 0$}}}{{\Longrightarrow}}0$ (36)
$\mathbb{F}_{\theta}^{2}$ $=\frac{3\sin^{2}\phi}{(1-|g|\cos\phi)(5+4|g|\cos\phi)}|g|^{2}\epsilon\stackrel{{\scriptstyle\text{$|g|\rightarrow 0$, $\phi\rightarrow\frac{\pi}{2}$}}}{{\Longrightarrow}}\frac{3}{5}|g|^{2}\epsilon$ (37)
Unfortunately, we don’t have a close form answer on the achievable Fisher information ratio $\mathbb{f}^{N}_{\theta}$ and $\mathbb{f}^{N}_{|g|}$ using their scheme, because the computational complexity of their protocols (And actualy, any protocols that involves $N$-copy SPE states and linear optical circuit $\mathcal{U}$) is essentially related to the computational complexity of Boson sampling[53]. However, from our observable, we conject that:
Conjecture 1.
linear optical implementation of $N$-copy SPE achieves optimal quantum telescopy asymptotically, however, the FI ratio $\mathbb{f}_{\mu}^{N}$ does not saturate the QFI ratio $\mathbb{h}^{N}$ and scales $\mathbb{f}_{\mu}^{N}\approx 1-\frac{2}{N+3}\stackrel{{\scriptstyle\text{Large $N$}}}{{\approx}}1-\frac{2}{N}$ for $\mu\in[|g|,\theta]$
One could plug in $\phi=0$ for $\mathbb{f}_{|g|}^{N}$ and $\phi=\pi/2$ for $\mathbb{f}_{\theta}^{N}$ to Eq. 34 to check our conjecture.
3 KLM and modified KLM schemes
As we have learned from the last section, the optimal ancilla state (using the least amount of photons) has the form: $|\psi\rangle_{a}=\sum_{n=0}^{N}f_{n}|n_{A},(N-n)_{B}\rangle$. From now on, we consider shared ancilla state with
$|\psi\rangle_{a}=\sum_{n=0}^{N}f_{n}|n_{A},(N-n)_{B}\rangle=\sum_{n=0}^{N}f_{n}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$ (38)
specifically, since they all promise a linear optical implementation
In this subsection, we will discuss schemes that are inspired by teleportation tricks used in linear optical quantum computation. [22, 32]. In the original paper by Knill, a $2N$-mode state of the following form is prepared
$|\psi\rangle^{\text{KLM}}_{a}=\frac{1}{\sqrt{N+1}}\sum\limits_{n=0}^{N}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}.$
We will detail the implementation of state teleportation using the shared entangled state of the above form later. But briefly speaking, in the original KLM teleportation[22], the teleportation trick could get the state teleported perfectly with $\frac{1}{N+1}$ chance fail (Happens when no photon arrived at telescope A or telescope B), we could also apply our Proposition 1 and obtain.
$\mathbb{h}[|\psi\rangle_{a}^{\text{KLM}}]=\frac{N}{N+1},$ (39)
which scales the same as the $N$-copy SPE, however, this bound is achievable using LOCC and linear optics.
A natural question arises: are there other shared entangled states, with different amplitudes $f_{n}$, that could achieve better quantum teleportation and, consequently, a higher quantum Fisher information ratio? Similar questions have been explored before, with answers quantified by the average fidelity of the teleported states [32]. The intuition behind these protocols is that by allowing a smaller chance of teleportation failure (less than $\frac{1}{N+1}$ as in the KLM scheme), one can accept a trade-off where each successful teleportation introduces a small error.
It is important to note that failed events are characterized by cases where no photon arrives at either telescope A or telescope B. A natural approach to shaping the amplitude $f_{n}$ is to reduce the values of $f_{0}$ and $f_{N+1}$. Inspired by this idea, we analyze in detail two modified KLM schemes: one where the intensity $|f_{n}|^{2}$ is shaped as a triangle, and another where the amplitude $|f_{n}|$ is shaped as a triangle.
In the first modification, we use
$|\psi\rangle^{\text{Itri}}_{a}=\sum\limits_{n=0}^{N}f^{\text{Itri}}_{n}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$
with $|f^{\text{Itri}}_{n}|^{2}\propto\frac{N}{2}-|\frac{N}{2}-n|$, i.e., the intensity of different terms shaped as a triangle. We assume $N$ as an even number for simplicity (the odd number case follows similarly), we have $|f^{\text{Itri}}_{n}|^{2}=\frac{4}{N^{2}}(\frac{N}{2}-|\frac{N}{2}-n|)$, which gives:
$\mathbb{h}[|\psi\rangle^{\text{Itri}}_{a}]=\frac{16}{N^{2}}\sum_{n=1}^{N/2}\frac{i(i-1)}{i+i-1}=\frac{16}{N^{2}}\sum_{n=1}^{N/2}(\frac{i}{2}-\frac{1}{4}-\frac{1}{4}\frac{1}{2i-1})=1-\frac{4}{N^{2}}\sum_{n=1}^{N/2}\frac{1}{2i-1}\stackrel{{\scriptstyle\text{Large {\hbox{N}}}}}{{\approx}}1-2\frac{\log(N)}{N^{2}},$ (40)
where we used the fact that $\sum_{n=1}^{N/2}\frac{1}{2i-1}\approx\frac{1}{2}(\log(2N)+\gamma)$.
Similarly, we could modify the amplitude even shaper such that we have
$|\psi\rangle^{\text{Atri}}_{a}=\sum\limits_{n=0}^{N}f^{\text{Atri}}_{n}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$
where $|f^{\text{Atri}}_{n}|\propto\frac{N}{2}-|\frac{N}{2}-n|$, i.e., the amplitude of different terms shaped as a triangle. Where again, take $N$ to be an even number for simplicity, we have $|f^{\text{Atri}}_{n}|^{2}=\frac{12}{N(N^{2}+2)}(\frac{N}{2}-|\frac{N}{2}-n|)^{2}$, which gives:
$\mathbb{h}[|\psi\rangle^{\text{Atri}}_{a}]=\frac{48}{N(N^{2}+2)}\sum_{n=1}^{N/2}\frac{i^{2}(i-1)^{2}}{i^{2}+(i-1)^{2}}=\frac{48}{N(N^{2}+2)}\sum_{n=1}^{N/2}(\frac{i^{2}}{2}-\frac{i}{2}-\frac{1}{4}+\frac{1}{4}\frac{1}{2i^{2}-2i+1})\stackrel{{\scriptstyle\text{Large {\hbox{N}}}}}{{\approx}}1-12\frac{1}{N^{2}},$ (41)
Which achieves a better QFI ratio asymptotically, and is only worse from the analytical upper bound by a constant factor. However, as we plot in Fig. 2 in the main text, for smaller $N$ the first modified-KLM scheme works better.
4 Optimal-KLM scheme
Now, we are going to show the best scheme that we have proposed, which is also quite close to the numerical optimal we could find. The shared entangled state we consider here takes the form
$|\psi\rangle^{\text{Sub}}_{a}=\sqrt{\frac{2}{N+2}}\sum\limits_{n=0}^{N}\sin(\frac{n+1}{N+2}\pi)|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B},$ (42)
which comes from the state we use to get analytical upper bound, and has been used in several other quantum information tasks [26, 51, 19]. Recall that, in lemma 4, we have shown that a state vector of the above form maximizes the term $\sum_{n=1}^{N}|f_{n}f_{n-1}|$, however, the quantify we want to optimize is the quantum Fisher information ratio:
$\mathbb{h}[\rho^{AB}_{a}]=\sum_{n=1}^{N}\frac{2|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}$
Plug $|f_{i}|=\sqrt{\frac{2}{N+2}}\sin(\frac{i+1}{N+2}\pi)$ in, we have:
$\mathbb{h}[\rho^{AB}_{a}]$ $=\frac{4}{N+2}\sum_{n=1}^{N}\frac{\sin^{2}(\frac{n+1}{N+2}\pi)\sin^{2}(\frac{n}{N+2}\pi)}{\sin^{2}(\frac{n+1}{N+2}\pi)+\sin^{2}(\frac{n}{N+2}\pi)}=\frac{1}{N+2}\sum_{n=1}^{N}\frac{[\cos(\frac{\pi}{N+2})-\cos(\frac{2n+1}{N+2}\pi)]^{2}}{1-\cos(\frac{\pi}{N+2})\cos(\frac{2n+1}{N+2}\pi)}$
$=\frac{\cos(\frac{\pi}{N+2})}{N+2}\sum_{n=1}^{N}[\cos(\frac{\pi}{N+2})-\cos(\frac{2n+1}{N+2}\pi)]-\frac{\sin^{2}(\frac{\pi}{N+2})}{N+2}\sum_{n=1}^{N}\frac{\cos(\frac{2n+1}{N+2}\pi)[\cos(\frac{1}{N+2}\pi)-\cos(\frac{2n+1}{N+2}\pi)]}{1-\cos(\frac{\pi}{N+2})\cos(\frac{2n+1}{N+2}\pi)}$ (43)
where the first term can be simplified as:
$\frac{\cos(\frac{\pi}{N+2})}{N+2}\sum_{n=1}^{N}[\cos(\frac{\pi}{N+2})-\cos(\frac{2n+1}{N+2}\pi)]=\cos^{2}(\frac{\pi}{N+2})$
For the second term, the coefficient $\frac{\sin^{2}(\frac{\pi}{N+2})}{N+2}=O(\frac{1}{N^{3}})$, and the summation is less than a negative constant.
Define that $f(n):=\frac{\cos(\frac{\pi}{N+2})-\cos(\frac{2n+1}{N+2}\pi)}{1-\cos(\frac{\pi}{N+2})\cos(\frac{2n+1}{N+2}\pi)}$, we have $f(n)\leq\cos(\frac{1}{N+2}\pi)$ for $\cos(\frac{2n+1}{N+2}\pi)\geq 0$ and $f(n)\geq\cos(\frac{1}{N+2}\pi)$ for $\cos(\frac{2n+1}{N+2}\pi)\leq 0$:
$\sum_{n=1}^{N}\cos(\frac{2n+1}{N+2}\pi)f(n)\leq\sum_{n=1}^{N}\cos(\frac{2n+1}{N+2}\pi)\cos(\frac{\pi}{N+2})\leq-\cos^{2}(\frac{\pi}{N+2})$
thus we conclude.
$\mathbb{h}[\rho^{AB}_{a}]\approx\cos^{2}(\frac{\pi}{N+2})\stackrel{{\scriptstyle\text{Large {\hbox{\langle N\rangle}}}}}{{\approx}}1-\frac{\pi^{2}}{N^{2}}$ (44)
5 Two-mode squeezed vacuum state
The two-mode squeezed vacuum state (TMSV) has served as the standard bipartite entangled state in continuous-variable systems [54, 55]. It was pointed out in the original quantum telescopy paper by Gottesman [10], and later in several other works [17, 18], that continuous-variable quantum teleportation [56, 57] using TMSV can facilitate a quantum telescopy. This seemingly contradicts the CV-protocol introduced in [17, 18], but aligns with the recent argument in [34], where TMSV is utilized, but CV-teleportation is not implemented.
A TMSQ state is written as
$|\psi\rangle^{\text{CV}}_{a}=\frac{1}{\cosh(r)}\sum\limits_{n=0}^{N}\tanh^{n}{r}|n\rangle_{A}|n\rangle_{B}$ (45)
where the squeezing parameter $r$ is related to the total average photon number $\langle N\rangle=2\sinh^{2}(r)$, so equivalently, one could also express it as:
$|\psi\rangle^{\text{CV}}_{a}=\sum\limits_{n=0}^{\infty}\frac{\sqrt{2\langle N\rangle^{n}}}{\sqrt{(2+\langle N\rangle)^{n+1}}}|n\rangle_{A}|n\rangle_{B}]$ (46)
At first glance, from Proposition 1, we have $\mathbb{h}[|\psi\rangle^{\text{CV}}_{a}]=0$ However, we aim to resolve this apparent contradiction.
In the standard continuous-variable quantum teleportation protocol [57], three steps are involved:
  • (1)
    Mixing the state (to be teleported) with one mode of the TMSV on a beam splitter.
  • (2)
    Applying a homodyne measurement to the joint state after mixing.
  • (3)
    Sending the measurement outcomes of the homodyne measurement $(x_{1},p_{2})$ (two quadrature at two outputs) is send to the other mode of TWSQ, where a displacement operator $\mathcal{D}(x_{1},p_{2})$ is implemented correspondingly.
It has been noted in the past that such a CV-teleportation can not been made if no extra resource is included[35, 36]. More precisely, an extra pair of correlated coherent state $|\alpha\rangle_{A}|\alpha\rangle_{B}$ need to be shared, such that Homodyne measurement at step (2) and displacement operator at step (3) can be done (note that this is different from the classical information send from A to B, but an extra pre-shared reference to ensure the Homodyne measurement and displacement operator are performed with the same phase reference).
With the extra classical correlated coherent states, one can group the two-mode squeezed vacuum state and shared coherent state and define
$|\psi\rangle^{\text{CV}}_{a}=\frac{1}{\cosh(r)}\sum\limits_{n=0}^{N}\tanh^{n}{r}|n\rangle_{A}|n\rangle_{B}|\alpha\rangle_{A}|\alpha\rangle_{B},$ (47)
which has non-vanishing Quantum Fisher information according to Proposition 1. To avoid the messy calculation because a perfect CV-teleportation protocol requires $|\alpha|\rightarrow\infty$, here we use results derived in the previous work[17]
$\mathbb{f}_{\theta}=\frac{2\epsilon^{2}|g|^{2}}{2y+\epsilon(2+\epsilon-\epsilon|g|^{2}+2y)}$ (48)
$\mathbb{f}_{|g|}=\frac{2\epsilon^{2}[-\epsilon(2+\epsilon)^{2}+\epsilon^{3}|g|^{4}-4(1+\epsilon)(2+\epsilon)y-4(2+\epsilon)y^{2}]}{[\epsilon(-1+|g|^{2})-2y][\epsilon(-2-\epsilon+\epsilon|g|^{2})-2(1+\epsilon)y][\epsilon^{2}(-1+|g|^{2})-4(1+y)-2\epsilon(2+y)]}$ (49)
Where $y$ is the squeezing parameter $y=2e^{-2r}$. Replacing $y=2(\sqrt{\frac{\langle N\rangle}{2}}+\sqrt{\frac{\langle N\rangle}{2}+1})^{-2}$, as $\langle N\rangle\rightarrow\infty$ we have, $y=\frac{1}{N}$ and thus:
$\mathbb{h}[|\psi\rangle^{\text{CV}}_{a}]\geq\mathbb{f}[|\psi\rangle^{\text{CV}}_{a}]\stackrel{{\scriptstyle\text{ {\hbox{N\gg\frac{1}{\epsilon}}}}}}{{\approx}}1-\frac{1}{\epsilon\langle N\rangle},$ (50)
Which is still worse than the optimal scheme we proposed in the above sections even if we have assume that we have an extra unbounded correlated coherent state $|\alpha\rangle_{A}|\alpha\rangle_{B}$.
6 Two particle entangled state, and NOON state
In this section, we attempt to use our framework, to answer questions that has been asked in the community quite a lot, that is, whether a more standrad two-photon two-mode Entangled state can be beneficial in quantum telescopy?
A two-photon entangled state is normally defined as :
$|\psi\rangle^{\text{TPE}}_{a}=\frac{1}{\sqrt{2}}(|10\rangle_{A}|01\rangle_{B}+|01\rangle_{A}|10\rangle_{B})$
This state represents a form of standard two-photon entangled states, also known as dual-rail encoding in quantum information theory [33] (while the single-particle entangled state is referred to as single-rail encoding). Nevertheless, from Proposition 1, this state yields zero quantum Fisher information.
The fact that the standard two-photon entangled state is useless in a quantum telescopy is not surprising, as it cannot be used as a resource to build shared phase references at all (which is defined as the PR resource in our main text). This feature has been highlighted in previous work [28] in a different context and is used to illustrate the distinction between nonlocal (entangled) resources and phase reference (PR) resources.
Before concluding the discussion on two-photon entangled states, we want to point out that there exists a straightforward way to activate the two-particle state: by simply ignoring the second mode on both telescopes, the state becomes identical to the single-particle entangled state used in Gottesman’s scheme, $|\psi\rangle^{\text{TPE}}_{a}\Rightarrow|\psi\rangle^{\text{SPE}}_{a}$. This preprocessing step can improve the QFI of the composite system in the presence of a superselection rule. However, since all the states being preprocessed into are still of the general form defined earlier, our upper bound on the QFI ratio remains universal, as established in Theorem 1.
Following the same spirit, now we can brief examine the usefulness of the so called NOON state, given as:
$|\psi\rangle^{\text{NOON}}=\frac{1}{\sqrt{2}}|N\rangle_{A}|0\rangle_{B}+|0\rangle_{A}|N\rangle_{B}$ (51)
We could quickly check that for $N\neq 1$, we have $\mathbb{h}[|\psi\rangle^{\text{NOON}}]=0$, similarly to what have observed in the two-photon entangled states. However, in this case, there exists no deterministic way of converting the NOON state to single-photon entangled states, however, probabilistic, it is still possible.
7 Correlated coherent states
Opposite to the example given in the previous section, where a nonlocal state can not be used to build shared phase references, here we give another of state that can be used as a PR resource, however, is not nonlocal thus leads to a gap between Fisher information and quantum fisher information.
The correlated coherent state has long known as a classical resources for building shared phase references, given as:
$|\psi\rangle^{\text{Coh}}_{a}=|\alpha\rangle_{A}|\alpha\rangle_{B}=\sum_{n,m=0}^{\infty}\frac{e^{-|\alpha|^{2}}\alpha^{n+m}}{\sqrt{n!m!}}|n\rangle_{A}|m\rangle_{B}$
It is not difficult to see that, for a coherent state, the quantum Fisher information $\mathbb{h}[|\psi\rangle^{\text{Coh}}_{a}$] is not vanishing with $f(n,m)=\frac{e^{-|\alpha|^{2}}\alpha^{n+m}}{\sqrt{n!m!}}$, and for large $|\alpha|^{2}$, one has numerically:
$\mathbb{h}[|\psi\rangle^{\text{Coh}}_{a}]\stackrel{{\scriptstyle\text{Large {\hbox{|\alpha|^2}}}}}{{\approx}}=1-\frac{1}{2|\alpha|^{2}}=1-\frac{1}{\langle N\rangle}$ (52)
where $\langle N\rangle=2|\alpha|^{2}$ is the total averaged photon used.
However, this doesn’t contradict the fact that coherent state/Homodyne measurement can only achieve a Fisher information scale $\epsilon^{2}$[11]. Instead, this is an example showing that Local operation and classical communication cannot always saturate the Quantum Fisher information, This indicates that, to ensure a shared ancillar state is useful for quantum telescopy, besides examining whether a shared state can be used to build a shared phase references, it is also necessary to check whether the state is entangled. Similar to what has already been proved in [11], we could conclude
Proposition 2.
If $[\rho_{a}^{AB}]^{T_{A}}>0$ (positive partial transpose), the FI ratio (instead of QFI ratio) will always go to 0.
Proof.
When an ancilla state $\rho_{A}$ is provided, the most general POVM that one could inplemented on the source can be written as convex combination of the one below.
$E(y)=\textrm{Tr}_{a}[U_{sa}^{A}(y)\otimes U_{sa}^{B}(y)\sum_{n,m=0}^{\infty}(P^{A}_{n}\otimes P^{B}_{m})\rho_{s}^{AB}\otimes\rho_{a}^{AB}(P^{A}_{n}\otimes P^{B}_{m})U_{sa}^{A\dagger}(y)\otimes U_{sa}^{B\dagger}(y)]$ (53)
Where $E(y)$ is the measurement operator with outcome $y$, and $U_{sa}^{A/B}(y)$ are local unitary at telescope $A$ and $B$
An immediate consequence of $[\rho_{a}^{AB}]^{T_{A}}>0$ is $E(y)^{T_{A}}>0$, and following what has been proved in [11], if a measurement operator is positive partial transpose, we have the Fisher information ratio.
$\mathbb{f}=O(\epsilon)\stackrel{{\scriptstyle\epsilon\rightarrow 0}}{{\Longrightarrow}}0$ (54)
Therefore, for a state with PR properties, i.e., $\mathbb{h}>0$, it is possible to find a gap between FI-ratio with all LOCC and QFI-ratio. The correlated coherent states will be a good example to that.
Appendix E Implementation with linear optical quantum teleportation
In this section, we will give some details on implementing several of our entanglement-assisted quantum telescopy protocols. The main of this linear optical implementation is brought from earlier work on linear optical quantum computation[22, 33], however, since those protocol are not given in too much detailed before in the previous work, and we are focusing on a different tasks focusing on maximizing the Fisher information, we will elaborate them in a little bit more detail.
Figure 4: Schematic of near-deterministic teleportation: (1) Quantum Fourier transformation $\mathcal{F}_{N+1}$ is implemented at telescope A; (2) Measurement outcome $n$: number of photons detected, and $\vec{s}$ arrangement of detection is sent to telescope B for phase correction; (3) A scanning interferometric measurement is performed as the direct interference scheme.
Here, we assume the source state is pure for simplicity with $|\psi\rangle_{s}^{(1)}=\frac{1}{\sqrt{2}}(|0\rangle_{A}|1\rangle_{B}+e^{i\theta}|1\rangle_{A}|0\rangle_{B})$, and the case for mixed source state follows straightforwardly by taking convex combination. Given a source state of the form of Eq. 38, we have the composite system (before applying superselection rule) as:
$|\psi\rangle^{AB}=\frac{1}{\sqrt{2}}(|0\rangle_{A}|1\rangle_{B}+e^{i\theta}|1\rangle_{A}|0\rangle_{B})\sum_{n=0}^{N}f_{n}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$ (55)
In the presence of superselection rule, we have the joint state $\rho^{AB}=\oplus_{k=0}^{N}\rho_{n}^{AB}$, where $\rho_{n}^{AB}=p_{n}|\psi\rangle\langle\psi|$ with
$p_{n}$ $=\frac{|f_{n}|^{2}}{2}\quad\quad\quad n=0,\quad N+1$ (56)
$p_{n}$ $=\frac{{|f_{n}|^{2}+|f_{n-1}^{2}|}}{2}\quad\quad\quad n=[1,\cdots,N]$ (57)
representing the normalized state where $n$-photon are detected and telescope $A$, which is given as:
$|\psi\rangle_{0}^{AB}$ $=|0\rangle^{N+1}_{A}|1\rangle_{B}^{N+1}\quad\quad\text{and}\quad\quad|\psi\rangle_{N+1}^{AB}=|1\rangle^{N+1}_{A}|0\rangle_{B}^{N+1}$
$|\psi\rangle_{n}^{AB}$ $=\frac{1}{\sqrt{|f_{n}|^{2}+|f_{n-1}|^{2}}}(f_{n}|0\rangle_{A}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|1\rangle_{B}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}+e^{i\theta}f_{n-1}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n+1}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n+1}_{B})$ (58)
The first step of the protocol consists of a quantum Fourier transformation on the $N+1$ mode at telescope A with $\mathcal{F}_{N+1}(a^{\dagger}_{p})=\sum_{q}w^{pq}a^{\dagger}_{q}=\sum_{q}\exp[2\pi i\frac{pq}{N+1}]a^{\dagger}_{q}$ for $p,q\in[0,\cdots,N]$, note that we are assuming the photons are indistinguishable either from the astronomical source of from the laboratory. For $n\in[1,\cdots,N-1]$ we have
$|\psi\rangle_{n}^{AB}$ $=\frac{1}{\sqrt{|f_{n}|^{2}+|f_{n-1}|^{2}}}\left(f_{n}\prod_{i=1}^{n}a^{\dagger}_{i}b^{\dagger}_{0}\prod_{j=n+1}^{N}b^{\dagger}_{j}+e^{i\theta}f_{n-1}\prod_{i=0}^{n-1}a^{\dagger}_{i}\prod_{j=n}^{N}b^{\dagger}_{j}\right)|\text{{0}}\rangle^{AB}$
$\mathcal{F}_{N+1}|\psi\rangle_{n}^{AB}$ $=\frac{1}{\sqrt{|f_{n}|^{2}+|f_{n-1}|^{2}}}\left(f_{n}\prod_{i=1}^{n}\sum\omega^{iq}a^{\dagger}_{q}b^{\dagger}_{0}\prod_{j=n+1}^{N}b^{\dagger}_{j}+e^{i\theta}f_{n-1}\prod_{i=0}^{n-1}\omega^{iq}a^{\dagger}_{q}\prod_{j=n}^{N}b^{\dagger}_{j}\right)|\text{{0}}\rangle^{AB}$ (59)
A certain arrangement $\vec{s}$ is a vector of length $N+1$, representing how many photons are detected at each output port labeled from $[0,\cdots,N]$. As $\vec{s}=[s_{0},\cdots,s_{N}]$ with $\sum_{i=0}^{N}s_{i}=n$. Since both terms in Eq. 59 has contribution to a certain arrangement, there is a phase differences to be fixed, more specifically, the for an arrangement $\vec{s}=[s_{0},\cdots,s_{N}]$ corresponds to a projection to $|\Psi(\vec{s})\rangle^{A}=\frac{1}{\sqrt{\prod_{i}s_{i}!}}\sum_{0}^{N}(a^{\dagger}_{i})^{s_{i}}|\text{{0}}\rangle^{A}$, we have:
$\prescript{A}{}{\langle\Psi(\vec{s})|}\mathcal{F}_{N+1}|\psi\rangle_{n}^{AB}$ $=\frac{1}{\sqrt{(|f_{n}|^{2}+|f_{n-1}|^{2})\prod_{n}s_{n}!}}\left(f_{n}c_{0}b^{\dagger}_{0}\prod_{j=k+1}^{N}b^{\dagger}_{j}+e^{i\theta}f_{n-1}c_{1}\prod_{j=k}^{N}b^{\dagger}_{j}\right)|\text{{0}}\rangle^{B}$ (60)
with $c_{0}=\sum_{\pi\in P^{0}_{n}}\prod_{i=1}^{n}\omega^{{d}_{i}(\vec{s})\pi(i)}\quad\quad c_{1}=\sum_{\pi\in P^{1}_{n}}\prod_{i=0}^{n-1}\omega^{{d}_{i}(\vec{s})\pi(i)}$ (61)
where $\pi$ is element in permutation group $P^{0}_{n}$ or $P^{1}_{n}$, which are permutation acting on the modes $[1,\cdots,n]$, or modes $[0,\cdots,n-1]$ (remember the 00th input mode is the source modes.) , and given an arrangement $\vec{s}$, a vector $\vec{d}$ is defined as a length-$n$ vector that specify each particle’s output port[53], i.e., $\vec{s}=[2,1,0]\Rightarrow\vec{d}(\vec{s})=[1,1,2]$, it is not difficult to see that,
$c_{0}=c_{1}\omega^{\sum_{i}d_{i}(\vec{s})}=c_{1}\exp[2\pi i\frac{\sum_{i}{d}_{i}(\vec{s})}{N+1}]$ (62)
Therefore, only a conditional phase gate $\mathcal{P}_{\phi(\vec{s})}$ with $\phi(\vec{s})=2\pi\frac{\sum_{i}{d}_{i}(\vec{s})}{N+1}$ need to introduce to correct the relatively phase, thus we could obtain phase correct states that can be coherently combined, which yields
$|\psi\rangle_{n}^{B}=\sum_{\vec{s}}\mathcal{P}_{\phi(\vec{s})}\prescript{A}{}{\langle\Psi(\vec{s})|}\mathcal{F}_{N+1}|\psi\rangle_{n}^{AB}=\frac{1}{\sqrt{|f_{n}|^{2}+|f_{n-1}|^{2}}}\left(f_{n}b^{\dagger}_{0}\prod_{j=n+1}^{N}b^{\dagger}_{j}+e^{i\theta}f_{n-1}\prod_{j=n}^{N}b^{\dagger}_{j}\right)|\text{{0}}\rangle^{B}$ (63)
And simply by ignoring all other modes but keep the 00-th and $n$-th modes at Bob’s side, with $n$ being the total number of photon detected at telescope A, we have
$|\psi\rangle_{n}^{B}=\frac{1}{\sqrt{|f_{n}|^{2}+|f_{n-1}|^{2}}}\left(f_{n}b^{\dagger}_{0}+e^{i\theta}f_{n-1}b^{\dagger}_{n}\right)|\text{{0}}\rangle^{B},$ (64)
by convexity, assuming a mixed source state $\rho_{s}^{(1)}$ of the form of Eq. 3 with $g=|g|e^{i\theta}$, we have:
$\rho_{k}^{B}=\frac{1}{|f_{n}|^{2}+|f_{n-1}|^{2}}\begin{pmatrix}|f_{n}|^{2}&f_{n}f^{*}_{n-1}g\\ f^{*}_{n}f_{n-1}g^{*}&|f_{n-1}|^{2}\end{pmatrix},\quad\quad\quad\text{Written in }\{|0\rangle_{B}|1_{n}\rangle_{B},|1\rangle_{B}|0_{n}\rangle_{B}\}\text{ basis.}$. (65)
Followed by a tilted beam splitter measurement characterized by projection to |±$|\pm\rangle_{n}\propto f_{n-1}|0\rangle_{B}|1_{n}\rangle_{B}+f_{n}e^{i\delta}|1_{n}\rangle_{B}|0\rangle_{B}$ with a tunable phase delay $\delta$, we have
$P^{n}_{\pm}$ $=p_{n}\langle\pm|\rho_{n}^{B}|\pm\rangle_{n}=\frac{|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}(1\pm|g|\cos(\theta-\delta))\quad n\in[1,\cdots,N]$ (66)
where $p_{n}$ is given in Eq. 57. One can hence, evaluate the Fisher information for the corresponding scheme as:
$\mathbb{F}_{|g|}$ $=\sum_{n,\pm}\frac{1}{P^{n}_{\pm}}\left(\frac{\partial P^{n}_{\pm}}{\partial|g|}\right)^{2}=\sum_{n=1}^{N}\frac{2|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}\frac{\cos(\theta-\delta)^{2}}{1-|g|^{2}\cos(\theta-\delta)^{2}}\epsilon+O(\epsilon^{2})\xrightarrow{\theta=\delta,\epsilon\ll 1}\sum_{n=1}^{N}\frac{2|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}\frac{\epsilon}{1-|g|^{2}}$ (67)
$\mathbb{F}_{|\theta|}$ $=\sum_{n,\pm}\frac{1}{P^{n}_{\pm}}\left(\frac{\partial P^{n}_{\pm}}{\partial|g|}\right)^{2}=\sum_{n=1}^{N}\frac{2|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}\frac{\sin(\theta-\delta)^{2}|g|^{2}}{1-|g|^{2}\cos(\theta-\delta)^{2}}\epsilon+O(\epsilon^{2})\xrightarrow{\theta=\delta+\frac{\pi}{2},\epsilon\ll 1}\sum_{n=1}^{N}\frac{2|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}}|g|^{2}\epsilon$ (68)
$\mathbb{f}$ $=\sum_{n=1}^{N}\frac{2|f_{n}|^{2}|f_{n-1}|^{2}}{|f_{n}|^{2}+|f_{n-1}|^{2}},$ (69)
which saturates the Quantum Fisher information we described in Proposition 1 for any state of the form:
$|\psi\rangle_{a}=\sum_{n=0}^{N}f_{n}|1\rangle^{\otimes n}_{A}|0\rangle^{\otimes N-n}_{A}|0\rangle^{\otimes n}_{B}|1\rangle^{\otimes N-n}_{B}$
Therefore, the KLM scheme, modified KLM schemes, and Optimal-KLM schemes we have introduced, are all implementable with local operation and classical communication with linear optics.